### v1

# On regions of mixed unitarity for semigroups of unital quantum channels

Abstract.
It is established that both discrete and continuous semigroups of unital quantum channels are eventually mixed unitary. However, after introducing the mixed unitary index of a unital quantum channel as the least time beyond which all subsequent powers of the channel are mixed unitary, we demonstrate that for any fixed finite dimension $d\geq 3$, there exists no universal upper bound for this index. Furthermore, for a continuous semigroup that is not mixed unitary for some $t>0$, we prove it remains non-mixed unitary for all times $t>0$, sufficiently close to the origin. Finally, a necessary and sufficient condition is derived for a quantum dynamical semigroup to be a convex combination of maps implemented solely by Weyl unitaries.
Key words and phrases:
Completely positive maps, quantum channels, semigroup, mixed unitary channels.
2020 Mathematics Subject Classification:
81P47, 47D06, 47D07, 47L90, 47L05
1. Introduction
In quantum information theory, a quantum channel (or simply a channel) serves as a fundamental mechanism for transmitting quantum information. Formally, it is defined as a trace-preserving, completely positive linear map from one matrix algebra to another.
For a natural number $d$, let $\mathbb{M}_{d}$ denote the algebra of all $d\times d$ complex matrices. In this article, we will be restricting ourselves to studying unital (doubly-stochastic) quantum channels from $\mathbb{M}_{d}$ to $\mathbb{M}_{d}$. A unitary quantum channel on $\mathbb{M}_{d}$ is a quantum channel which is given by $X\mapsto U^{*}XU$, for all $X\in\mathbb{M}_{d}$ and for some unitary $U\in\mathbb{M}_{d}$. Unitary quantum channels are extreme in the convex set of unital quantum channels on $\mathbb{M}_{d}$. A quantum channel is said to be a mixed unitary if it is in the convex hull of the set of all unitary quantum channels, or equivalently (due to finite dimensionality) it is a convex combination of unitary channels. That is, a quantum channel $\Phi:\mathbb{M}_{d}\to\mathbb{M}_{d}$ is a mixed unitary if there exist $\{\lambda_{j}\}_{j=1}^{\ell}\subseteq[0,1]$ and unitaries $\{U_{j}\}_{j=1}^{\ell}\subseteq\mathbb{M}_{d}$ for some $l\in\mathbb{N}$ such that $\sum_{j=1}^{\ell}\lambda_{j}=1$ and
$\Phi(X)=\sum_{i=1}^{\ell}\lambda_{i}U_{i}^{*}XU_{i},\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \forall X\in\mathbb{M}_{d}.$ (1.1)
The set of all mixed unitary channels on $\mathbb{M}_{d}$ forms a compact convex subset of all unital quantum channels.
Mixed unitary channels play a significant role in quantum information theory for several key reasons. They provide a rich class of examples that capture key properties of general quantum operations [Ros08]. By virtue of their structural simplicity, they remain highly useful for theoretical analysis, often proving more tractable than arbitrary quantum channels. These channels emerge naturally in multiple contexts, including majorization frameworks based on Hermitian operators [AU82], as well as in quantum cryptographic protocols involving state encryption[AMTB00, HDPW04, AS04].
Taking inspiration from the classical Birkhoff-von Neumann theorem, which realizes all doubly stochastic maps as convex combinations of permutations, one may suspect that perhaps every unital quantum channel is mixed unitary. A well-known result [LaSt93, Theorem 1] establishes this for $d=2$. However, this property fails for $d\geq 3$ as first demonstrated by Tregub [Tre86] and Kümmerer-Maassen [KuMa87]. See also [LaSt93, BPS93]. Additionally, determining whether a given unital channel is mixed unitary or not is known to be NP-hard [LWat20].
Fix a dimension $d\in\mathbb{N}$ and consider the completely depolarizing channel on $\mathbb{M}_{d}:$
$\delta_{d}(X):=\frac{\mbox{trace}(X)}{d}I_{d},\penalty 10000\ \penalty 10000\ X\in\mathbb{M}_{d},$
A famous theorem of Watrous [Wat09], says that there is a ball of positive radius around $\delta_{d}$ such that all unital channels in it are mixed unitary. As a consequence, any sequence of unital channels converging to the completely depolarizing channel eventually consists of only mixed unitary channels. This was taken note of in the context of powers of primitive unital channels in Section 4 of [HRS20]. They define mixed unitary time (or index) of a unital quantum channel as a threshold $k$ such that all subsequent powers beyond which are mixed unitary. Moreover they show that such a threshold exists for primitive unital channels and obtain an upper bound for it.
In a more recent work, Kribs et al., [KJRM23] extend this idea very significantly. Note that the completely depolarizing channel can be viewed as the unique conditional expectation map from $\mathbb{M}_{d}$ to the $C^{*}$-algebra $\mathbb{C}I_{d}$ of scalars. Fix a unital $C^{*}$-subalgebra $\mathcal{A}$ of $\mathbb{M}_{d}$ and let $\mathbf{E}_{\mathcal{A}}$ be the unique trace-preserving conditional expectation to it. It is a unital quantum channel. In [KJRM23] it is observed that unital quantum channels sufficiently close to $\mathbf{E}_{\mathcal{A}}$ and fixing elements of $\mathcal{A}$ are mixed unitary. Armed with this result they show that if $\Phi$ is a unital quantum channel on $\mathbb{M}_{d}$, then there exists some $k\in\mathbb{N}$ such that $\Phi^{k}$ is mixed unitary [KJRM23, Theorem 22]. However, $\Phi^{k}$ fixes elements of an algebra $\mathcal{A}$, does not imply that $\Phi^{k+1}$ does the same. For this reason, they couldn’t show that $\{\Phi^{n}\}_{n\geq 0}$ has finite mixed unitary time. We overcome this difficulty in Theorem 3.9, by making use of the notion of conditional automorphism on the peripheral eigenspace of $\Phi$.
This raises the natural question as to for fixed $d$, whether there exists a bound $N_{d}$ such that mixed unitary indices of all unital quantum channels on $\mathbb{M}_{d}$ are bounded by $N_{d}$. We answer this question in the negative in Theorem 4.9. We get it as a consequence of our analysis of continuous ( one parameter) semigroups of unital quantum channels.
A deep study of continuous semigroups of mixed unitary channels was carried out long time ago by Kümmerer and Maassen in [KuMa87]. Their goal was to understand Markov dilations of such semigroups. In the process they characterized generators of quantum dynamical semigroups which remain mixed unitary for all times. We obtain a different characterization of such generators (Theorem 5.2). We show that even in continuous time semigroups of unital channels eventually become mixed unitary. In other words, they all have finite mixed unitary index (Theorem 4.7).
We organize the paper as follows: In Section 2, we recall the necessary definitions and basic results required in the subsequent sections, and fix some notations. Section 3, investigates the mixed unitarity for a discrete semigroup $\{\Phi^{n}\}_{n\in\mathbb{N}}$ generated by an unital quantum channel $\Phi$ on $\mathbb{M}_{d}$. We first establish that every trace-preserving *-automorphism on a $C^{*}$-subalgebra $\mathcal{A}$ of $\mathbb{M}_{d}$ is necessarily of the form $\mathrm{Ad}_{U}$ for some $U\in\mathbb{M}_{d}$ (see Theorem 3.6). As a consequence of this Theorem, we prove that any unital quantum channel on $\mathbb{M}_{d}$ is asymptotically a conditional *-automorphism (Theorem 3.8). This leads us to one of the main results, namely every discrete semigroup $\{\Phi^{n}\}_{n\in\mathbb{N}}$ of unital quantum channels on $\mathbb{M}_{d}$ is eventually mixed unitary (Theorem 3.9).
In section 4, we extend the analysis to continuous one parameter semigroups $\{\Phi_{t}\}_{t\geq 0}$ of unital quantum channels on $\mathbb{M}_{d}$. An analogous theorem is established, proving these semigroups are also eventually mixed unitary (Theorem 4.7). We also prove that there is no universal upper bound of the mixed unitary index for a fixed dimension (Theorem 4.9). That is, there is no $N\geq 1$ such that $\Phi^{N}$ is a mixed unitary channel, for all unital channels $\Phi$ on $\mathbb{M}_{d}$.
In Section 5, we examine the local behavior of a continuous semigroup of a unital quantum channels $\{\Phi_{t}\}_{t\geq 0}$ near $t=0$. First, a characterization theorem is established for such a semigroup on $\mathbb{M}_{d}$ to be mixed unitary for all times (Theorem 5.2). Using this Theorem, we prove that every continuous one parameter semigroup $\{\Phi_{t}\}_{t\geq 0}$ of unital quantum channels on $\mathbb{M}_{d}$ is either mixed unitary for all times or there is a time $t_{0}$ such that $\Phi_{t}$ is never mixed unitary for all times $t\in(0,t_{0})$ (Theorem 5.4). In the last section, we discuss mixed Weyl-unitary channels on $\mathbb{M}_{d}$. We provide a complete structure theorem (Theorem 6.4) for a one-parameter semigroup $\{\Phi_{t}\}_{t\geq 0}$ of unital quantum channels on $\mathbb{M}_{d}$ to be mixed Weyl-unitary for all times. We also introduce the notion of $G$-mixed unitary channel on $\mathbb{M}_{d}$ for any subgroup $G$ of $\mathbb{U}(d)$. For any closed subgroup $G$, we have an analogous theorem (see Theorem 6.8) for a continuous one parameter semigroup $\{\Phi_{t}\}_{t\geq 0}$ of unital quantum channels on $\mathbb{M}_{d}$ to be $G$-mixed unitary for all times.
2. Background
2.1. Notations
Let $\mathbb{C}^{d}$ denote the $d$-dimensional complex Hilbert space with the inner product linear in the second variable and conjugate linear in the first. The set $\{e_{i}:1\leq i\leq d\}$ is the standard orthonormal basis of $\mathbb{C}^{d}$. Let $\mathbb{M}_{d_{1}\times d_{2}}$ be the algebra of complex matrices of size $d_{1}\times d_{2}$, we use $\mathbb{M}_{d}$ to refer $\mathbb{M}_{d\times d}$. The identity matrix in $\mathbb{M}_{d}$ is denoted as $I=I_{d}$. Additionally, we let $\operatorname{\textsf{T}}=\operatorname{\textsf{T}}_{d}:\mathbb{M}_{d}\to\mathbb{M}_{d}$ be the transpose map (w.r.t. the standard basis), $\operatorname{\textsf{tr}}=\operatorname{\textsf{tr}}_{d}:\mathbb{M}_{d}\to\mathbb{C}$ be the trace map and $\operatorname{\textsf{id}}=\operatorname{\textsf{id}}_{d}:\mathbb{M}_{d}\to\mathbb{M}_{d}$ be the identity map. By writing $A=[a_{ij}]\in\mathbb{M}_{d_{1}\times d_{2}}$, we mean that $A$ is a complex matrix of size $d_{1}\times d_{2}$, with $(i,j)^{th}$ entry equal to $a_{ij}$, for all indices satisfying $1\leq i\leq d_{1}$ and $1\leq j\leq d_{2}$. If $A=[a_{ij}]\in\mathbb{M}_{d_{1}\times d_{2}}$, the adjoint of $A$, denoted as $A^{*}$, is a matrix $A^{*}=[b_{ij}]\in\mathbb{M}_{d_{2}\times d_{1}}$, where each entry $b_{ij}$ is given by $\overline{a_{ji}}$ for all $1\leq i\leq d_{2}$ and $1\leq j\leq d_{1}$. A matrix $A\in\mathbb{M}_{d}$ is said to be Hermitian if $A=A^{*}$, and a Hermitian matrix $A$ is said to be positive (semi-definite) if $\langle x,Ax\rangle\geq 0$ for all $x\in\mathbb{C}^{d}$. The collection of all positive matrices in $\mathbb{M}_{d}$ is denoted by $\mathbb{M}_{d}^{+}$. If $A\in\mathbb{M}_{d}^{+}$, then we write $A\geq 0$. A matrix $U\in\mathbb{M}_{d}$ is said to be unitary if $U^{*}U=UU^{*}=I$, and denotes the set of all $d\times d$ unitary matrices by $\mathbb{U}(d)$.
For every $x\in\mathbb{C}^{d_{1}}$ and $y\in\mathbb{C}^{d_{2}}$, define the rank one linear map $|{x}\rangle\langle{y}|:\mathbb{C}^{d_{2}}\to\mathbb{C}^{d_{1}}$ by $|{x}\rangle\langle{y}|(z):=\langle y,z\rangle x$, for all $z\in\mathbb{C}^{d_{2}}$. It is straightforward to verify that $A|{x}\rangle\langle{y}|B=|{Ax}\rangle\langle{B^{*}y}|$ holds for any $x\in\mathbb{C}^{d_{1}}$, $y\in\mathbb{C}^{d_{2}}$, and matrices $A\in\mathbb{M}_{d_{1}},B\in\mathbb{M}_{d_{2}}$. Additionally, it is easy to see that the matrix $|{x}\rangle\langle{x}|$ is positive for all $x\in\mathbb{C}^{d}$.
Given the linear maps $\Phi_{1}:\mathbb{M}_{d_{1}}\to\mathbb{M}_{d_{2}}$ and $\Phi_{2}:\mathbb{M}_{d_{3}}\to\mathbb{M}_{d_{4}}$, the tensor product map $\Phi_{1}\otimes\Phi_{2}:\mathbb{M}_{d_{1}}\otimes\mathbb{M}_{d_{3}}\to\mathbb{M}_{d_{2}}\otimes\mathbb{M}_{d_{4}}$ is the unique map determined by $\Phi_{1}\otimes\Phi_{2}(A\otimes B)=\Phi_{1}(A)\otimes\Phi_{2}(B)$ for all elements $A\in\mathbb{M}_{d_{1}}$ and $B\in\mathbb{M}_{d_{3}}$.
Definition 2.1.
A linear map $\Phi:\mathbb{M}_{d_{1}}\to\mathbb{M}_{d_{2}}$ is said to be
  1. (i)
    Hermitian preserving if $\Phi(X^{*})=\Phi(X)^{*}$ for all $X\in\mathbb{M}_{d_{1}}$;
  2. (ii)
    positive if $\Phi(\mathbb{M}_{d_{1}}^{+})\subseteq\mathbb{M}_{d_{2}}^{+}$;
  3. (iii)
    completely positive (CP) if the map $\operatorname{\textsf{id}}_{k}\otimes\Phi:\mathbb{M}_{k}\otimes\mathbb{M}_{d_{1}}\to\mathbb{M}_{k}\otimes\mathbb{M}_{d_{2}}$ is positive, for all $k\geq 1$;
  4. (iv)
    unital if $\Phi(I_{d_{1}})=I_{d_{2}}$;
  5. (v)
    trace-preserving if $\operatorname{\textsf{tr}}(\Phi(X))=\operatorname{\textsf{tr}}(X)$ for all $X\in\mathbb{M}_{d_{1}}$;
  6. (vi)
    unital quantum channel if $\Phi$ is a unital and trace-preserving CP map.
Let $\langle X,Y\rangle:=\operatorname{\textsf{tr}}(X^{*}Y)$ be the Hilbert-Schmidt inner product on $\mathbb{M}_{d}$. Then the dual of a linear map $\Phi:\mathbb{M}_{d_{1}}\to\mathbb{M}_{d_{2}}$ is the unique linear map $\Phi^{*}:\mathbb{M}_{d_{2}}\to\mathbb{M}_{d_{1}}$ such that $\operatorname{\textsf{tr}}({\Phi(X)}^{*}Y)=\operatorname{\textsf{tr}}(X^{*}\Phi^{*}(Y))$ for all $X\in\mathbb{M}_{d_{1}}$ and $Y\in\mathbb{M}_{d_{2}}$. Denote the space of all linear maps from $\mathbb{M}_{d_{1}}$ to $\mathbb{M}_{d_{2}}$ by $\mathscr{B}({\mathbb{M}_{d_{1}},\mathbb{M}_{d_{2}}})$. In quantum information theory, the Choi-Jamiołkowski isomorphism [Cho75, Jam72], commonly known as channel-state duality, illustrates the connection between quantum channels and quantum states.
Theorem 2.2 (Choi-Jamiołkowski [Cho75, Jam72]).
Let $\mathscr{J}$ be the linear map from $\mathscr{B}({\mathbb{M}_{d_{1}},\mathbb{M}_{d_{2}}})\to\mathbb{M}_{d_{1}}\otimes\mathbb{M}_{d_{2}}$ defined by
$\Phi\mapsto(\operatorname{\textsf{id}}_{d_{1}}\otimes\Phi)(|{\Omega_{d_{1}}}\rangle\langle{\Omega_{d_{1}}}|),$
where $\Omega_{d_{1}}=\frac{1}{\sqrt{d_{1}}}\sum_{j=1}^{d_{1}}e_{j}\otimes e_{j}$. Then we have the following:
  1. (i)
    $\mathscr{J}$ is a vector space isomorphism.
  2. (ii)
    $\Phi$ is Hermitian preserving if and only if $\mathscr{J}(\Phi)$ is Hermitian.
  3. (iii)
    $\Phi$ is positive map if and only if $\mathscr{J}(\Phi)$ is positive on simple tensor vectors, that is, $\langle x\otimes y,\mathscr{J}(\Phi)x\otimes y\rangle\geq 0$ for all $x\in\mathbb{C}^{d_{1}}$ and $y\in\mathbb{C}^{d_{2}}$.
  4. (iv)
    $\Phi$ is CP if and only if $\mathscr{J}(\Phi)$ is positive.
  5. (v)
    $\Phi$ is unital if and only if $(\operatorname{\textsf{tr}}\otimes\operatorname{\textsf{id}}_{d_{2}})(\mathscr{J}(\Phi))=\frac{I_{d_{2}}}{d_{1}}$.
  6. (vi)
    $\Phi$ is trace-preserving if and only if $(\operatorname{\textsf{id}}_{d_{1}}\otimes\operatorname{\textsf{tr}})(\mathscr{J}(\Phi))=\frac{I_{d_{1}}}{d_{1}}$.
The matrix $\mathscr{J}(\Phi)$ is known as the Choi matrix of $\Phi.$
Corollary 2.3 (Choi-Kraus [Cho75, Kra71]).
Let $\Phi:\mathbb{M}_{d_{1}}\to\mathbb{M}_{d_{2}}$ be a linear map. Then $\Phi$ is CP if and only if there exists $\{V_{j}\}_{j=1}^{n}\subseteq\mathbb{M}_{d_{1}\times d_{2}}$ such that
$\Phi=\sum_{j=1}^{n}\mathrm{Ad}_{V_{j}},$ (2.1)
where $\mathrm{Ad}_{V}(X):=V^{*}XV$ for all $X\in\mathbb{M}_{d_{1}}$.
The representation given by (2.1) is known as the Kraus decomposition of the map $\Phi$, and the operators $V_{i}$’s are called the Kraus operators. The minimal number of Kraus operators required to represent $\Phi$ is called the Choi-rank of $\Phi$, and it is equal to the rank of the Choi-matrix $\mathscr{J}(\Phi)$. Note that the Kraus decomposition of $\Phi$ need not be unique. However, if $\Phi$ is expressed in two different Kraus decomposition $\sum_{j=1}^{n}\mathrm{Ad}_{V_{j}}$ and $\sum_{i=1}^{m}\mathrm{Ad}_{W_{i}}$, then $\operatorname{\textsf{span}}\{V_{j}:1\leq j\leq n\}=\operatorname{\textsf{span}}\{W_{i}:1\leq i\leq m\}$.
Definition 2.4.
An unital quantum channel $\Phi:\mathbb{M}_{d}\to\mathbb{M}_{d}$ is said to be mixed unitary if there exist $\{\lambda_{j}\}_{j=1}^{\ell}\subseteq[0,1]$ and unitaries $\{U_{j}\}_{j=1}^{\ell}\subseteq\mathbb{U}(d)$ such that $\sum_{j=1}^{\ell}\lambda_{j}=1$ and
$\Phi(X)=\sum_{i=1}^{\ell}\lambda_{i}U_{i}^{*}XU_{i},\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \forall X\in\mathbb{M}_{d}.$
The mixed unitary rank of $\Phi$ is the minimal $\ell$ for which such decomposition exists. It is immediate that, Choi-rank of $\Phi$ is less than or equal to the mixed unitary rank of $\Phi$. In [Bus06], it is proved that the mixed unitary rank of $\Phi$ is at most the square of the Choi-rank of $\Phi$, which is further less than or equal to $d^{4}$. For more details on mixed unitary rank, see [CKLi22].
It is well-known that every unital quantum channel on $\mathbb{M}_{2}$ is mixed unitary [Wat18, KuMa87]. This is not true in general for $d\geq 3$. For ready reference we have the following example.
Example 2.5.
[Wat18, Example 4.3] (Holevo-Werner channel) Let $\Phi:\mathbb{M}_{3}\to\mathbb{M}_{3}$ be the unital quantum channel given by
$\Phi(X):=\frac{tr(X)I_{3}-X^{T}}{2},\quad\penalty 10000\ \penalty 10000\ \forall X\in\mathbb{M}_{3}.$
Then $\Phi$ is not a mixed unitary.
Let $\mathscr{H}(d)$ and $\mathscr{MU}(d)$ be the set of all Hermitian-preserving linear maps on $\mathbb{M}_{d}$ and the collection of all mixed unitary-channels on $\mathbb{M}_{d}$, respectively. The Choi-Jamiołkowski isomorphism $\mathscr{J}$ (c.f. 2.2) establishes a bijective correspondence between $\mathscr{B}({\mathbb{M}_{d},\mathbb{M}_{d}})$ and $\mathbb{M}_{d}\otimes\mathbb{M}_{d}$. Define an inner product on $\mathscr{B}({\mathbb{M}_{d}}):=\mathscr{B}({\mathbb{M}_{d},\mathbb{M}_{d}})$ by
$\langle\Phi,\Psi\rangle:=\operatorname{\textsf{tr}}\big(\mathscr{J}(\Phi)^{*}\mathscr{J}(\Psi)\big).$ (2.2)
Then, $\mathscr{J}$ becomes an isometric isomorphism between two Hilbert spaces. Furthermore, we have the following:
  1. (i)
    The set $\mathscr{H}(d)$ is isometrically isomorphic to the subset of $\mathbb{M}_{d}\otimes\mathbb{M}_{d}$ consisting of Hermitian matrices, denoted by $(\mathbb{M}_{d}\otimes\mathbb{M}_{d})_{Her}$;
  2. (ii)
    The image of $\mathscr{MU}(d)$ under the Choi-Jamiołkowski isomorphism, denoted by $\mathscr{J}(\mathscr{MU}(d))$, is equal to the convex hull of Choi matrices corresponding to unitary channels. i.e., $\mathscr{J}(\mathscr{MU}(d))=\text{co}\{\mathscr{J}({\mathrm{Ad}_{U}}):U\in\mathbb{U}(d)\}$ and this subset is contained within the set of all positive matrices $\rho\in\mathbb{M}_{d}\otimes\mathbb{M}_{d}$ such that $(\operatorname{\textsf{id}}\otimes\operatorname{\textsf{tr}})(\rho)=(\operatorname{\textsf{tr}}\otimes\operatorname{\textsf{id}})(\rho)=\frac{1}{d}I_{d}$.
Recall that a subset $\mathcal{C}$ of a real topological vector space $\mathcal{V}$ is a convex cone if it is a convex set and closed under non-negative scalar multiplication. Let $\mathcal{V}^{{}^{\prime}}$ denote the set of all continuous functionals on $\mathcal{V}$.
Definition 2.6.
Let $\mathcal{C}$ be any subset of $\mathcal{V}$. The dual of $\mathcal{C}$, denoted and defined as
$\mathcal{C}^{\circ}:=\{f\in\mathcal{V}^{{}^{\prime}}:f(c)\geq 0,\forall c\in\mathcal{C}\}$ (2.3)
Note that $\mathcal{C}^{\circ}$ is a closed convex cone. Furthermore, if $\mathcal{V}$ is the locally convex Hausdorff space, then the double dual of a set $\mathcal{C}$ is equal to the closed convex cone generated by $\mathcal{C}$. The following lemma is known in the literature, but we include the details for completeness.
Lemma 2.7.
Consider $\mathscr{MU}(d)$ as a subset of the real vector space $\mathscr{H}(d)$. Then the following holds:
  1. (i)
    $\mathscr{MU}(d)$ is a compact convex set and is closed under compositions.
  2. (ii)
    $\mathscr{MU}(d)^{\circ}=\{\Gamma\in\mathscr{H}(d):\langle\Gamma,\mathrm{Ad}_{U}\rangle\geq 0\penalty 10000\ \penalty 10000\ \forall U\in\mathbb{U}(d)\}$.
  3. (iii)
    Suppose $\Phi\in\mathscr{H}(d)$ is unital and trace-preserving. Then $\Phi\in\mathscr{MU}(d)$ if and only if $\langle\Gamma,\Phi\rangle\geq 0$ for all unital and trace-preserving $\Gamma\in\mathscr{MU}(d)^{\circ}$.
  4. (iv)
    The interior of the set $\mathscr{MU}(d)$ is empty.
Proof.
$(i)$ Since the set set of unitary channels is compact and $\mathscr{MU}(d)$ is its convex hull it is also compact. Clearly it is closed under compositions.
$(ii)$ Note that, any linear functional $f$ on $\mathscr{H}(d)$ is of the form $f(\Phi)=\operatorname{\textsf{tr}}\big(\mathscr{J}(\Gamma)\mathscr{J}(\Phi)\big)=\langle\Gamma,\Phi\rangle$, for some $\Gamma\in\mathscr{H}(d)$. Therefore, from the definition of the dual cone,
$\mathscr{MU}(d)^{\circ}$ $=\{f\in(\mathscr{H}(d))^{{}^{\prime}}:f(\Phi)\geq 0,\penalty 10000\ \penalty 10000\ \forall\Phi\in\mathscr{MU}(d)\}$
$=\{\Gamma\in\mathscr{H}(d):\langle\Gamma,\Phi\rangle\geq 0,\penalty 10000\ \penalty 10000\ \forall\Phi\in\mathscr{MU}(d)\}$
$=\{\Gamma\in\mathscr{H}(d):\langle\Gamma,\mathrm{Ad}_{U}\rangle\geq 0,\penalty 10000\ \penalty 10000\ \forall U\in\mathbb{U}(d)\}.$
$(iii)$ Assume that $\Phi$ is mixed unitary, that is $\Phi=\sum_{j=1}^{\ell}\lambda_{j}\mathrm{Ad}_{U_{j}}$ for some $U_{j}\in\mathbb{U}(d)$ and $\lambda_{j}\in[0,1]$, then for any unital, trace-preserving $\Gamma\in\mathscr{MU}(d)^{\circ}$, we have
$\langle\Gamma,\Phi\rangle$ $=\langle\Gamma,\sum_{j=1}^{\ell}\lambda_{j}\mathrm{Ad}_{U_{j}}\rangle=\sum_{j=1}^{\ell}\lambda_{j}\langle\Gamma,\mathrm{Ad}_{U_{j}}\rangle\geq 0.$
Conversely, assume that $\langle\Gamma,\Phi\rangle\geq 0$ for all unital and trace-preserving $\Gamma\in\mathscr{MU}(d)^{\circ}$. Let $\delta_{d}(X)=\frac{1}{d}\operatorname{\textsf{tr}}(X)I_{d},\penalty 10000\ \penalty 10000\ \forall X\in\mathbb{M}_{d}$, and consider the sets,
$\mathcal{M}$ $=\{\Psi-\delta_{d}:\Psi\in\mathscr{MU}(d)\},$
$\mathcal{N}:$ $=\{\Psi-\delta_{d}:\Psi\in\mathscr{H}(d)\text{ is unital and trace-preserving}\}.$
Since $\mathscr{MU}(d)$ is a closed convex set, we have $\mathcal{M}$ is a closed convex subset of a real vector space $\mathcal{N}$. Suppose $\Phi\notin\mathscr{MU}(d)$, then $\mathscr{J}(\Phi-\delta_{d})\in\mathscr{J}(\mathcal{N})\setminus\mathscr{J}(\mathcal{M})$. Note that any linear functional on $\mathscr{J}(\mathcal{N})$ is of the form $Z\mapsto\operatorname{\textsf{tr}}(YZ)$ for some $Y\in\mathscr{J}(\mathcal{N})$. Therefore, by the Hahn-Banach separation theorem  [^1] Theorem [Ze12, Theorem 1. C]: Let $M$ be a closed non-empty convex subset of a normed linear space $\mathcal{X}$ over $\mathbb{R}$. Let $u_{0}\in\mathcal{X}$ but not in $M$, then there exists a continuous linear function $f$ on $\mathcal{X}$ such that $f(u)\leq 1<f(u_{0})$ for all $u\in M$. there exist $Y\in\mathscr{J}(\mathcal{N})$ such that
$\operatorname{\textsf{tr}}(YZ)\leq 1<\operatorname{\textsf{tr}}(Y\mathscr{J}(\Phi-\delta_{d})),\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \forall Z\in\mathscr{J}(\mathcal{M}).$
Take $\Gamma:=\mathscr{J}^{-1}(W)$, where $W=\frac{1}{d^{2}}(I_{d^{2}}-Y)$. Then $\Gamma\in\mathscr{H}(d)$. Now as $(\operatorname{\textsf{id}}\otimes\operatorname{\textsf{tr}})(W)=\frac{I_{d}}{d}=(\operatorname{\textsf{tr}}\otimes\operatorname{\textsf{id}})(W)$, we have $\Gamma$ is unital and trace-preserving. Furthermore, for every unitary $U\in\mathbb{U}(d)$,
$\langle\Gamma,\mathrm{Ad}_{U}\rangle=\operatorname{\textsf{tr}}(\mathscr{J}(\Gamma)\mathscr{J}(\mathrm{Ad}_{U}))$ $=\operatorname{\textsf{tr}}(W\mathscr{J}(\mathrm{Ad}_{U}))$
$=\frac{1}{d^{2}}\{\operatorname{\textsf{tr}}(\mathscr{J}(\mathrm{Ad}_{U}))-\operatorname{\textsf{tr}}(Y\mathscr{J}(\mathrm{Ad}_{U}))\}$
$=\frac{1}{d^{2}}\{1-\operatorname{\textsf{tr}}(Y\mathscr{J}(\mathrm{Ad}_{U}))+\frac{1}{d^{2}}\operatorname{\textsf{tr}}(Y)\}\qquad(\because\operatorname{\textsf{tr}}(Y)=0,$ $\operatorname{\textsf{tr}}(\mathscr{J}(\mathrm{Ad}_{U}))=1)$
$=\frac{1}{d^{2}}\{1-\operatorname{\textsf{tr}}\left(Y(\mathscr{J}(\mathrm{Ad}_{U})-\frac{1}{d^{2}}I_{d^{2}})\right)\}$
$=\frac{1}{d^{2}}\{1-\operatorname{\textsf{tr}}\left(Y\big(\mathscr{J}(\mathrm{Ad}_{U}-\delta_{d})\big)\right)\}$
$\geq 0$
This shows that $\Gamma\in\mathscr{MU}(d)^{\circ}$. But,
$\langle\Gamma,\Phi\rangle=\operatorname{\textsf{tr}}(\mathscr{J}(\Gamma)\mathscr{J}(\Phi))=\operatorname{\textsf{tr}}(W\mathscr{J}(\Phi))$ $=\frac{1}{d^{2}}\{\operatorname{\textsf{tr}}(\mathscr{J}(\Phi))-\operatorname{\textsf{tr}}(Y\mathscr{J}(\Phi))\}$
$=\frac{1}{d^{2}}\{1-\operatorname{\textsf{tr}}(Y\mathscr{J}(\Phi))+\frac{1}{d^{2}}\operatorname{\textsf{tr}}(Y)\}$
$=\frac{1}{d^{2}}(1-\operatorname{\textsf{tr}}(Y\mathscr{J}(\Phi-\delta_{d})))<0.$
This contradicts our assumption.
$(iv)$ Suppose the interior of $\mathscr{MU}(d)$ is non-empty. i.e., there exists a $\Phi_{0}\in\mathscr{MU}(d)$ and $r>0$ such that $B(\Phi_{0},r):=\{\Psi\in\mathscr{H}(d):\left\lVert\Psi-\Phi_{0}\right\rVert_{2}<r\}$ [^2] Since any two norms are equivalent in finite dimensional space, so we can choose the ball in Hilbert-Schmidt norm. is a subset of $\mathscr{MU}(d)$. Here $\left\lVert\Psi\right\rVert_{2}:=\operatorname{\textsf{tr}}\big(\mathscr{J}(\Psi)^{*}\mathscr{J}(\Psi)\big)$. Observe that $\mathscr{J}(B(\Phi_{0},r)-\delta_{d})\subseteq\mathcal{X}$, where $\mathcal{X}=\{\rho\in(\mathbb{M}_{d}\otimes\mathbb{M}_{d})_{Her}:\operatorname{\textsf{id}}\otimes\operatorname{\textsf{tr}}(\rho)=\operatorname{\textsf{tr}}\otimes\operatorname{\textsf{id}}(\rho)=0\}$ and $\delta_{d}=\frac{\operatorname{\textsf{tr}}(.)}{d}I_{d}$. Since $\mathscr{J}$ is an isometry and the translation of an open set is open, we deduce that $\mathcal{X}$ contains an open set. Therefore, $\mathcal{X}=(\mathbb{M}_{d}\otimes\mathbb{M}_{d})_{Her}$, which is not possible. ∎
3. Discrete semigroups of mixed unitary channels
Example 2.5 shows that for $d\geq 3$, not every unital quantum channel $\Phi$ on $\mathbb{M}_{d}$ is a mixed unitary. However, it is proven recently that the finite evolution of the unital quantum channel would lead to a mixed unitary at some point. More explicitly Theorem 21 of [KJRM23] shows that if $\Phi:\mathbb{M}_{d}\to\mathbb{M}_{d}$ is a unital quantum channel, then there exists some $k\in\mathbb{N}$ such that $\Phi^{k}$ is a mixed unitary channel. This suggests the following definition.
Definition 3.1.
Let $\{\Phi_{k}\}_{k\in\mathbb{N}}$ be a sequence of CP maps on $\mathbb{M}_{d}.$ (i) It is said to be frequently mixed unitary if it has a subsequence of mixed unitary channels. (ii) It is said to be eventually mixed unitary if there exists $k\in\mathbb{N}$ such that $\Phi_{n}$ is mixed unitary for $n\geq k.$
Clearly, if $\Phi^{k}$ is a mixed unitary so is $\Phi^{mk}$ for all $m\geq 1.$ Hence the main result of [KJRM23] is telling us that for any unital quantum channel $\Phi$, $\{\Phi^{n}\}_{n\in\mathbb{N}}$ is frequently mixed unitary. In this article (see Theorem 3.9), we make the stronger assertion that it is eventually mixed unitary. This was known in some special cases, such as primitive (meaning that it is irreducible and has a trivial peripheral spectrum) channels ([HRS20, KJRM23]) and here we have it in general. The main tool we need is a result from [KJRM23] which says that quantum channels fixing elements of a subalgebra $\mathcal{A}$ and sufficiently close to the conditional expectation map on to $\mathcal{A}$ are mixed unitary. We begin with relevant definitions and results.
Definition 3.2.
Let $\mathcal{A}$ be a unital $C^{*}$-subalgebra of $\mathbb{M}_{d}$. A CP map $\mathbf{E}:\mathbb{M}_{d}\to\mathbb{M}_{d}$ with $\mathbf{E}(\mathbb{M}_{d})=\mathcal{A}$ is said to be a conditional expectation onto $\mathcal{A}$ if $\mathbf{E}(A)=A,\forall A\in\mathcal{A}$.
Note that a conditional expectation map $\mathbf{E}$ as above is of norm 1 as it is a unital CP map and it is also a projection as it is the identity map on its range. It is also known that $\mathbf{E}(AXB)=A\mathbf{E}(X)B$ for all $A,B\in\mathcal{A}$ and $X\in\mathbb{M}_{d}.$ See [Ume53, Ra06].
Conditional expectation onto a $C^{*}$ subalgebra need not be unique. However, there is only one conditional expectation that preserves the trace, denote it by $\mathbf{E}_{\mathcal{A}}$. In fact, it is equal to the orthogonal projection of $\mathbb{M}_{d}$ onto $\mathcal{A}$ w.r.t. the Hilbert-Schmidt inner product on $\mathbb{M}_{d}$. In [BhBha25] one can see a concrete way of writing $\mathbf{E}_{\mathcal{A}}$ as a mixed unitary channel. One of the fundamental quantum channels widely discussed in the literature is the completely depolarizing channel $\delta_{d}$:
$\delta_{d}(X)=\frac{1}{d}\operatorname{\textsf{tr}}(X)I_{d},\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \forall X\in\mathbb{M}_{d}.$ (3.1)
Note that this is the unique trace-preserving conditional expectation onto the algebra $\mathbb{C}I_{d}$. The following Proposition is due to John Watrous[Wat09]. It says that all unital quantum channels sufficiently close to the depolarizing channel are mixed unitary.
Proposition 3.3.
[Wat09, Corollary 2] Fix $d\in\mathbb{N}.$ There exists $r>0$ depending on $d$ such that, every unital, quantum channel $\Phi$ on $\mathbb{M}_{d}$ with $\left\lVert\Phi-\delta_{d}\right\rVert<r$, is mixed unitary.
In 2024, David Kribs et al. [KJRM23] extended this result with the depolarising channel replaced by conditional expectation maps on arbitrary unital $C^{*}$-subalgebras of $\mathbb{M}_{d}$. For a given a unital $C^{*}$-subalgebra $\mathcal{A}$ of $\mathbb{M}_{d}$, we let $\mathcal{F}(\mathcal{A})$ to be the set of all unital quantum channels which leave elements of $\mathcal{A}$ fixed. That is,
$\mathcal{F}(\mathcal{A}):=\{\Phi:\mathbb{M}_{d}\to\mathbb{M}_{d}:\Phi\mbox{ is unital quantum channel such that }\Phi(A)=A,\penalty 10000\ \penalty 10000\ \forall A\in\mathcal{A}\}.$
It is easy to see that $\mathcal{F}(\mathcal{A})$ is closed under the compositions. Furthermore, it is closed under adjoints also. This follows from the fact that the fixed point space of $\Phi$ is equal to the commutant of the set $\{V_{j}:1\leq j\leq\ell\}$, where $\{V_{j}\}_{j=1}^{\ell}$ is any set of Kraus operators of $\Phi$ (See [Kri03, Theorem 2.1]).
The following result has been observed and used in [KJRM23]. However, it is not explicitly stated or proved. The result is essential for us and so we provide the details.
Theorem 3.4.
[KJRM23, Theorem 4, Lemma 12] Let $\mathcal{A}$ be a unital $C^{*}$-subalgebra of $\mathbb{M}_{d}$ and let $\mathbf{E}_{\mathcal{A}}$ be the unique trace-preserving conditional expectation onto $\mathcal{A}$. Then there exists $r_{\mathcal{A}}>0$ such that, if $\Phi\in\mathcal{F}(\mathcal{A})$ with $\left\lVert\Phi-\mathbf{E}_{\mathcal{A}}\right\rVert\leq r_{\mathcal{A}}$, then $\Phi$ is mixed unitary.
Proof.
Lemma 12 of [KJRM23] proves that $\mathbf{E}_{\mathcal{A}}$ is mixed unitary. Theorem 4 of the same article shows that there exists a $p\in(0,1)$ such that $p\Phi+(1-p)\mathbf{E}_{\mathcal{A}}$ is mixed unitary, for all $\Phi\in\mathcal{F}(\mathcal{A})$. Now, for a given $\Phi\in\mathcal{F}(\mathcal{A})$, define $\Phi_{p}:=\mathbf{E}_{\mathcal{A}}+\frac{1}{p}(\Phi-\mathbf{E}_{\mathcal{A}})$. Clearly, $\Phi_{p}$ fixes elements of $\mathcal{A}$. Suppose $\Phi_{p}$ is CP, then $\Phi_{p}\in\mathcal{F}(\mathcal{A})$, and hence $p\Phi_{p}+(1-p)\mathbf{E}_{\mathcal{A}}=\Phi$ is mixed unitary. Therefore, it suffices to prove that there exists $r_{\mathcal{A}}\geq 0$ such that $\left\lVert\Phi-\mathbf{E}_{\mathcal{A}}\right\rVert\leq r_{\mathcal{A}}$, then $\Phi_{p}$ is CP. For this, without loss of generality assume that $\mathcal{A}=\bigoplus_{k=1}^{\ell}(I_{m_{k}}\otimes\mathbb{M}_{n_{k}})$, where
$I_{m_{k}}\otimes\mathbb{M}_{n_{k}}:=\{I_{m_{k}}\otimes X\in\mathbb{M}_{m_{k}n_{k}}:X\in\mathbb{M}_{n_{k}}\}$
and $n_{k},m_{k}\in\mathbb{N}$ such that $\sum_{k=1}^{\ell}m_{k}n_{k}=d$. For computational simplicity, we will only prove the case when $\ell=2$, the general case follows similarly. Let $\mathcal{H}=\mathcal{H}_{1}\oplus\mathcal{H}_{2},$ where $H_{k}=\mathbb{C}^{m_{k}}\otimes\mathbb{C}^{n_{k}}$, for $k=1,2$ and $\bigcup_{k=1}^{2}\{e_{r}^{(k)}\otimes f_{s}^{(k)}:1\leq r\leq m_{k},1\leq s\leq n_{k}\}$ be the standard orthonormal basis of $\mathcal{H}$. Then, the associated matrix units of $\mathscr{B}({\mathcal{H}})$ are $\bigcup_{j,k=1}^{2}\{E_{rr^{\prime}}^{(j,k)}\otimes F_{ss^{\prime}}^{(j,k)}:1\leq r\leq m_{j},1\leq r^{\prime}\leq m_{k},1\leq s\leq n_{j},1\leq s^{\prime}\leq n_{k}\}$, where $E_{rr^{\prime}}^{(j,k)}=|{e_{r}^{(j)}}\rangle\langle{e_{r^{\prime}}^{(k)}}|$ and $F_{ss^{\prime}}^{(j,k)}=|{f_{s}^{(j)}}\rangle\langle{f_{s^{\prime}}^{(k)}}|$. Write any $A\in\mathscr{B}({\mathcal{H}})$ in the block form as $A=\begin{bmatrix}A_{11}&A_{12}\\ A_{21}&A_{22}\end{bmatrix}$, with $A_{jk}\in\mathscr{B}({\mathcal{H}_{k},\mathcal{H}_{j}})$, for $j,k=1,2$. Note that, the map $\mathbf{E}_{\mathcal{A}}$ is given by
$\mathbf{E}_{\mathcal{A}}\big(\begin{bmatrix}A_{11}&A_{12}\\ A_{21}&A_{22}\end{bmatrix})=\begin{bmatrix}\frac{I_{m_{1}}}{m_{1}}\otimes(\operatorname{\textsf{tr}}\otimes\operatorname{\textsf{id}})(A_{11})&0\\ 0&\frac{I_{m_{2}}}{m_{2}}\otimes(\operatorname{\textsf{tr}}\otimes\operatorname{\textsf{id}})(A_{22})\end{bmatrix}$.
Let $\{K_{j}\}_{j=1}^{N}$ be a set of Kraus operators of $\Phi$. Since $\Phi\in\mathcal{F}(\mathcal{A})$, we get that
$K_{j}=\begin{bmatrix}K_{j1}\otimes I_{n_{1}}&0\\ 0&K_{j2}\otimes I_{n_{2}}\end{bmatrix},\penalty 10000\ \penalty 10000\ \forall 1\leq j\leq N,$
with $K_{ji}\in\mathbb{M}_{m_{i}}$ such that $\sum_{j=1}^{N}K_{ji}^{*}K_{ji}=\sum_{j=1}^{N}K_{ji}K_{ji}^{*}=I_{m_{i}}$, for $i=1,2$. Therefore, the map $\Phi$ is given by
$\Phi\big(\begin{bmatrix}A_{11}&A_{12}\\ A_{21}&A_{22}\end{bmatrix}\big)=\begin{bmatrix}\sum_{j=1}^{N}(K_{j1}^{*}\otimes I_{n_{1}})A_{11}(K_{j1}\otimes I_{n_{1}})&\sum_{j=1}^{N}(K_{j1}^{*}\otimes I_{n_{1}})A_{12}(K_{j2}\otimes I_{n_{2}})\\ \sum_{j=1}^{N}(K_{j2}^{*}\otimes I_{n_{2}}A_{21}(K_{j1}\otimes I_{n_{1}})&\sum_{j=1}^{N}(K_{j2}\otimes I_{n_{2}})A_{22}(K_{j2}\otimes I_{n_{2}})\end{bmatrix}$.
and the Choi-matrix of $\Phi$ is given as
$\mathscr{J}(\Phi)$ $=\frac{1}{d}\Big(\sum_{j,k=1}^{2}\sum_{r,r^{\prime}=1}^{m_{j},m_{k}}\sum_{s,s^{\prime}=1}^{n_{j},n_{k}}E_{rr^{\prime}}^{(j,k)}\otimes F_{ss^{\prime}}^{(j,k)}\otimes\Phi\big(E_{rr^{\prime}}^{(j,k)}\otimes F_{ss^{\prime}}^{(j,k)}\big)\Big)$
$=\frac{1}{d}\begin{bmatrix}\sum_{r,r^{\prime}=1}^{m_{1}}\sum_{s,s^{\prime}=1}^{n_{1}}e_{rr^{\prime}}^{(1,1)}\otimes f_{ss^{\prime}}^{(1,1)}\otimes\sum_{j=1}^{N}(K_{j1}^{*}e_{rr^{\prime}}^{(1,1)}K_{j1})\otimes f_{ss^{\prime}}^{(1,1)}&\sum_{r,r^{\prime}=1}^{m_{1},m_{2}}\sum_{s,s^{\prime}=1}^{n_{1},n_{2}}e_{rr^{\prime}}^{(1,2)}\otimes f_{ss^{\prime}}^{(1,2)}\otimes\sum_{j=1}^{N}(K_{j1}^{*}e_{rr^{\prime}}^{(1,2)}K_{j2})\otimes f_{ss^{\prime}}^{(1,2)}\\ \sum_{r,r^{\prime}=1}^{m_{2},m_{1}}\sum_{s,s^{\prime}=1}^{n_{2},n_{1}}e_{rr^{\prime}}^{(2,1)}\otimes f_{ss^{\prime}}^{(2,1)}\otimes\sum_{j=1}^{N}(K_{j2}^{*}e_{rr^{\prime}}^{(2,1)}K_{j1})\otimes f_{ss^{\prime}}^{(2,1)}&\sum_{r,r^{\prime}=1}^{m_{2}}\sum_{s,s^{\prime}=1}^{n_{2}}e_{rr^{\prime}}^{(2,2)}\otimes f_{ss^{\prime}}^{(2,2)}\otimes\sum_{j=1}^{N}(K_{j2}^{*}e_{rr^{\prime}}^{(2,2)}K_{j2})\otimes f_{ss^{\prime}}^{(2,2)}\end{bmatrix}$,
$e_{rr^{\prime}}^{(j,k)},f_{ss^{\prime}}^{(j,k)}$ are the standard matrix units of $\mathbb{M}_{m_{j}\times m_{k}},\mathbb{M}_{n_{j}\times n_{k}}$, respectively. Note that if $\mathcal{K}=\mathcal{K}_{1}\oplus\mathcal{K}_{2}$, then the matrix units $e_{rr^{\prime}}$ in $\mathscr{B}({\mathcal{K}_{1}})$ is treated as $E_{rr^{\prime}}=\begin{bmatrix}e_{rr^{\prime}}&0\\ 0&0\end{bmatrix}$ in $\mathscr{B}({\mathcal{K}})$. Now, by swapping the second and third tensor products in each block entry, the Choi-matrix is equal to
$\frac{1}{d}\begin{bmatrix}A\otimes|{f^{(1)}}\rangle\langle{f^{(1)}}|&B\otimes|{f^{(1)}}\rangle\langle{f^{(2)}}|\\ C\otimes|{f^{(2)}}\rangle\langle{f^{(1)}}|&D\otimes|{f^{(2)}}\rangle\langle{f^{(2)}}|\end{bmatrix}$,
where
$A:$ $=\sum_{r,r^{\prime}=1}^{m_{1}}e_{rr^{\prime}}^{(1,1)}\otimes\sum_{j=1}^{N}(K_{j1}^{*}e_{rr^{\prime}}^{(1,1)}K_{j1}),\qquad B:=\sum_{r,r^{\prime}=1}^{m_{1},m_{2}}e_{rr^{\prime}}^{(1,2)}\otimes\sum_{j=1}^{N}(K_{j1}^{*}e_{rr^{\prime}}^{(1,2)}K_{j2})$
$C:$ $=\sum_{r,r^{\prime}=1}^{m_{2},m_{1}}e_{rr^{\prime}}^{(2,1)}\otimes\sum_{j=1}^{N}(K_{j2}^{*}e_{rr^{\prime}}^{(2,1)}K_{j1}),\qquad D:=\sum_{r,r^{\prime}=1}^{m_{2}}e_{rr^{\prime}}^{(2,2)}\otimes\sum_{j=1}^{N}(K_{j2}^{*}e_{rr^{\prime}}^{(2,2)}K_{j2}),$
and $f^{(j)}=\sum_{s}f_{s}^{(j)}\otimes f_{s}^{(j)}$, for $j=1,2$. Note that $\frac{1}{m_{1}+m_{2}}\begin{bmatrix}A&B\\ C&D\end{bmatrix}$ is the Choi-matrix (upto permutation equivalence) of the unital quantum channel $\Psi:\mathbb{M}_{m_{1}+m_{2}}\to\mathbb{M}_{m_{1}+m_{2}}$ given by
$\Psi(\begin{bmatrix}X_{11}&X_{12}\\ X_{21}&X_{22}\end{bmatrix})=\sum_{j=1}^{N}\begin{bmatrix}K_{j1}^{*}&0\\ 0&K_{j2}^{*}\end{bmatrix}\begin{bmatrix}X_{11}&X_{12}\\ X_{21}&X_{22}\end{bmatrix}\begin{bmatrix}K_{j1}&0\\ 0&K_{j2}\end{bmatrix}$.
Thus, by letting $\Psi^{\prime}:=(m_{1}+m_{2})\Psi$, the Choi-matrix (upto permutation equivalence) of $\Phi$ is equal to the tensorial Schur product (see [VsSu15]) of $\frac{1}{d}\mathscr{J}(\Psi^{\prime})$ and $|{f}\rangle\langle{f}|$. That is,
$\frac{1}{d}\mathscr{J}(\Psi^{\prime})\circ^{\otimes}|{f}\rangle\langle{f}|,$
where $f=f^{(1)}\oplus f^{(2)}$. It is a principal submatrix of $\frac{1}{d}\mathscr{J}(\Psi^{\prime})\otimes|{f}\rangle\langle{f}|$. Using the similar argument, we also have that the Choi-matrix (upto permutation equivalence) of $\mathbf{E}_{\mathcal{A}}$ is equal to
$\frac{1}{d}\begin{bmatrix}\frac{I_{m_{1}}\otimes I_{m_{1}}}{m_{1}}&0\\ 0&\frac{I_{m_{2}}\otimes I_{m_{2}}}{m_{2}}\end{bmatrix}\circ^{\otimes}|{f}\rangle\langle{f}|.$
Therefore, the Choi-matrix (upto permutation equivalence) of $\Phi_{p}$ is $\frac{1}{d}\big(Z+\frac{1}{p}(\mathscr{J}(\Psi^{\prime})-Z)\big)\circ^{\otimes}|{f}\rangle\langle{f}|,$ where $Z=\begin{bmatrix}\frac{I_{m_{1}}\otimes I_{m_{1}}}{m_{1}}&0\\ 0&\frac{I_{m_{2}}\otimes I_{m_{2}}}{m_{2}}\end{bmatrix}$. Now, as $Z$ is strictly positive, taking $r$ as the smallest positive eigenvalue of $Z$, $Z+\Delta$ is positive, whenever $\Delta=\Delta^{*}$ with $\left\lVert\Delta\right\rVert\leq r$. Since considering a principal submatrix can only reduce the norm, we get $\left\lVert\mathscr{J}(\Phi)-\mathscr{J}(\mathbf{E}_{\mathcal{A}})\right\rVert\geq\frac{1}{d}\left\lVert\mathscr{J}(\Psi^{\prime})-Z\right\rVert$. Set $r_{\mathcal{A}}=\frac{pr}{d}$, then whenever $\Phi\in\mathcal{F}(\mathcal{A})$ with $\left\lVert\mathscr{J}(\Phi)-\mathscr{J}(\mathbf{E}_{\mathcal{A}})\right\rVert\leq r_{\mathcal{A}}$ we have
$\frac{1}{p}\left\lVert\mathscr{J}(\Psi^{\prime})-Z\right\rVert=\frac{d}{p}\left\lVert\mathscr{J}(\Phi)-\mathscr{J}(\mathbf{E}_{\mathcal{A}})\right\rVert\leq r.$
Hence $\big(Z+\frac{1}{p}(\mathscr{J}(\Psi^{\prime})-Z)\big)\geq 0$. By [VsSu15, Proposition 1.3], $\frac{1}{d}\big(Z+\frac{1}{p}(\mathscr{J}(\Psi^{\prime})-Z)\big)\circ^{\otimes}|{f}\rangle\langle{f}|$ is positive. Further, this is equivalent to the map $\Phi_{p}$ being CP. This completes the proof. ∎
Definition 3.5.
A linear map $\Phi:\mathbb{M}_{d}\to\mathbb{M}_{d}$ said to be a conditional *-automorphism if $\Phi=\mathbf{E}_{\mathcal{A}}\circ\Pi$, where $\mathbf{E}_{\mathcal{A}}$ is a trace-preserving conditional expectation onto some $C^{*}$-subalgebra $\mathcal{A}$ of $\mathbb{M}_{d}$ and $\Pi$ is a *-automorphism on $\mathbb{M}_{d}$ which commutes with $\mathbf{E}_{\mathcal{A}}$.
In general an automorphsim of a subalgebra $\mathcal{A}$ of $M_{d}$ need not be implemented by a unitary. For instance, take $\mathcal{A}=\{a(E_{1}1+E_{22})+bE_{33}:a,b\in\mathbb{C}\}\subseteq\mathbb{M}_{3}$, where $E_{ij}$ denotes matrix units and consider the automorphism $a(E_{11}+E_{22})+bE_{33}\mapsto b(E_{11}+E_{22})+aE_{33}.$ However, there are no such issues if the automorphism preserves the trace as the following theorem shows.
Theorem 3.6.
Let $\mathcal{A}$ be a unital $C^{*}$-subalgebra of $\mathbb{M}_{d}$. Let $\Psi$ be a trace-preserving *-automorphism on $\mathcal{A}$, then there exists a unitary $U\in\mathbb{M}_{d}$ such that $\Psi=\mathrm{Ad}_{U}$.
Proof.
Without loss of generality assume that $\mathcal{A}=\oplus_{k=1}^{\ell}(I_{m_{k}}\otimes\mathbb{M}_{n_{k}})$, where
$I_{m_{k}}\otimes\mathbb{M}_{n_{k}}:=\{I_{m_{k}}\otimes X\in\mathbb{M}_{m_{k}n_{k}}:X\in\mathbb{M}_{n_{k}}\}$
and $n_{k},m_{k}\in\mathbb{N}$ such that $\sum_{k=1}^{\ell}n_{k}m_{k}=d$. (Note that $m_{k},n_{k}$ can be repeating.) Each summand
$J_{i}:=\bigoplus_{k=1}^{\ell}\delta_{ik}(I_{m_{k}}\otimes\mathbb{M}_{n_{k}}),$
is a minimal two-sided ideal of $\mathcal{A}$, and these are the only minimal two-sided ideals of $\mathcal{A}$. For any $1\leq k\leq\ell$ and $m\in\{m_{1},m_{2},\cdots,m_{\ell}\}$, let
$B_{k}(m):=\{i:J_{i}=\bigoplus_{k=1}^{\ell}\delta_{ik}(I_{m_{k}}\otimes\mathbb{M}_{n_{k}})\text{with }n_{i}=n_{k},m_{i}=m\}.$
[…]
$\Psi\left(\bigoplus_{r=1}^{\ell}\delta_{ir}(I_{m_{r}}\otimes X)\right)$ $=\bigoplus_{r=1}^{\ell}\delta_{\sigma_{k}(i)r}(I_{m_{r}}\otimes U_{\sigma_{k}(i)}^{*}XU_{\sigma_{k}(i)}),\penalty 10000\ \penalty 10000\ \forall X\in\mathbb{M}_{n_{k}}.$
Let $U_{k}=\bigoplus_{r=1}^{\ell}(I_{m_{r}}\otimes W_{r}),$ where $W_{r}=U_{\sigma_{k}(r)}$ if $r\in B_{k}(m)$, otherwise identity matrix in $\mathbb{M}_{n_{r}}$, and $P_{k}$ in $\mathbb{M}_{d}$ be the permutation matrix that reorders the blocks in $B_{k}(m)$ according to $\sigma_{k}$. That is,
$P_{k}^{*}\left(\sum_{i\in B_{k}(m)}\bigoplus_{r=1}^{\ell}\delta_{ir}(I_{m_{r}}\otimes X_{i})\right)P_{k}=\sum_{i\in B_{k}(m)}\bigoplus_{r=1}^{\ell}\delta_{\sigma_{k}(i)r}(I_{m_{r}}\otimes X_{i}),\penalty 10000\ \penalty 10000\ \penalty 10000\ \forall X_{i}\in\mathbb{M}_{n_{k}}.$
[…]
$\Psi\left(\sum_{i\in B_{k}(m)}\bigoplus_{r=1}^{\ell}\delta_{ir}(I_{m_{r}}\otimes X_{i})\right)$ $=\sum_{i\in B_{k}(m)}\bigoplus_{r=1}^{\ell}\delta_{\sigma_{k}(i)r}(I_{m_{k}}\otimes U_{\sigma_{k}(i)}^{*}X_{i}U_{\sigma_{k}(i)})$
$=P_{k}^{*}\left(\sum_{i\in B_{k}(m)}\bigoplus_{r=1}^{\ell}\delta_{ir}(I_{m_{k}}\otimes U_{\sigma_{k}(i)}^{*}X_{i}U_{\sigma_{k}(i)})\right)P_{k}$
$=P_{k}^{*}U_{k}^{*}\left(\sum_{i\in B_{k}(m)}\bigoplus_{r=1}^{\ell}\delta_{ir}(I_{m_{r}}\otimes X_{i})\right)U_{k}P_{k},$
for all $X_{i}\in\mathbb{M}_{n_{k}}$. Now, let $n_{1},n_{2},\cdots n_{q}$ be distinct numbers of the set $\{n_{1},n_{2},\cdots,n_{\ell}\}$ and $m_{1},m_{2}\cdots,m_{s}$ be the distinct numbers of the set $\{m_{1},m_{2},\cdots m_{\ell}\}$ . Then $\mathcal{A}=\sum_{k=1}^{q}\sum_{j=1}^{s}\mathcal{A}_{k}(m_{j})$, where $\mathcal{A}_{k}(m_{j}):=\sum_{i\in B_{k}(m_{j})}J_{i}$. Furthermore, for every $(k,j)\neq(k^{\prime},j^{\prime})$, $\mathcal{A}_{k}(m_{j})\mathcal{A}_{k^{\prime}}(m_{j^{\prime}})=0=\mathcal{A}_{k^{\prime}}(m_{j^{\prime}})\mathcal{A}_{k}(m_{j})$. Consequently, one can choose a global permutation matrix $P\in\mathbb{M}_{d}$ and a block diagonal unitary $V\in\mathbb{M}_{d}$ such that $\Psi(Z)=P^{*}V^{*}ZVP$, for all $Z\in\mathcal{A}$. Taking $U=VP$, completes the proof. ∎
It is to be noted that generally the unitary $U$ of this theorem is not an element of the algebra $\mathcal{A}.$ However, $\mathrm{Ad}_{U}$ leaves $\mathcal{A}$ and $\mathcal{A}^{\perp}$ invariant. There is extensive literature on the asymptotics of CP maps and quantum channels. We will follow mostly [BhSa23]. We cite [AFK25] for recent work and the citations in it for further information.
Definition 3.7.
Let $\Phi:\mathbb{M}_{d}\to\mathbb{M}_{d}$ be a linear map.
  1. (i)
    The multiplicative domain of $\Phi$, is defined as
    $\mathscr{M}_{\Phi}:=\{X\in\mathbb{M}_{d}:\Phi(XY)=\Phi(X)\Phi(Y)\mbox{ and }\Phi(YX)=\Phi(Y)\Phi(X)\quad\penalty 10000\ \forall Y\in\mathbb{M}_{d}\}.$
    It is known that for a unital CP map $\Phi$, we have
    $\mathscr{M}_{\Phi}=\{X\in\mathbb{M}_{d}:\Phi(X^{*}X)=\Phi(X^{*})\Phi(X)\mbox{ and }\Phi(XX^{*})=\Phi(X)\Phi(X^{*})\}.$
  2. (ii)
    The peripheral spectrum of $\Phi$ be the set of all elements in the spectrum $\sigma(\Phi)$ whose absolute value is $1$. i.e., $\{\lambda\in\sigma(\Phi):\left\lvert\lambda\right\rvert=1\}$.
  3. (iii)
    The peripheral space of $\Phi$ is defined as
    $\mathcal{P}(\Phi):=\operatorname{\textsf{span}}\{X:\Phi(X)=\lambda X,\left\lvert\lambda\right\rvert=1\}.$
It is known that for a unital CP map $\Phi$, we have
$\mathscr{M}_{\Phi}=\{X\in\mathbb{M}_{d}:\Phi(X^{*}X)=\Phi(X^{*})\Phi(X)\mbox{ and }\Phi(XX^{*})=\Phi(X)\Phi(X^{*})\}.$
Recall that for any CP map $\Phi$ on $\mathbb{M}_{d}$, we have $\Phi(X^{*})\Phi(X)\leq\left\lVert\Phi(I)\right\rVert\Phi(X^{*}X)$, for all $X\in\mathbb{M}_{d}$ (This inequality is known as the Kadison-Schwarz inequality). As a consequence of this inequality, we get that if $\Phi$ is a unital quantum channel, then the peripheral space $\mathcal{P}(\Phi)\subseteq\mathscr{M}_{\Phi}$.
Theorem 3.8.
Let $\Phi$ be a unital quantum channel on $\mathbb{M}_{d}.$ Then $\{\Phi^{n}\}_{n\in\mathbb{N}}$ is asymptotically a conditional *-automorphism. i.e., there exists a conditional *-automorphishm $\alpha$ on $\mathbb{M}_{d}$ such that
$\lim_{n\to\infty}\left\lVert\Phi^{n}-\alpha^{n}\right\rVert=0.$
Also there exists an automorphism $\gamma$ of $\mathbb{M}_{d}$, such that $(\gamma\circ\Phi)^{n}$ converges to the conditional expectation map onto the peripheral space of $\Phi$ as $n$ tends to $\infty.$
Proof.
Let $\mathcal{P}(\Phi)$ be the peripheral space of $\Phi$. That is, $\mathcal{P}(\Phi)=\operatorname{\textsf{span}}\{X:\Phi(X)=\lambda X,\left\lvert\lambda\right\rvert=1\}$. Then, the assumption that $\Phi$ is trace-preserving ensures that [BhSa23], $\mathcal{P}(\Phi)$ is a $C^{*}$-sub algebra of $\mathbb{M}_{d}.$ Let $\mathbf{E}_{\mathcal{P}}$ be the trace-preserving conditional expectation onto $\mathcal{P}(\Phi)$. Considering $\mathbb{M}_{d}$ as a Hilbert space with Hilbert-Schmidt inner product, $\mathbf{E}_{\mathcal{P}}$ is just the orthogonal projection on to the subspace $\mathcal{P}(\Phi).$ Moreover we have the following ([BhSa23], Theorem 3.1 and 3.4):
  1. (i)
    $\mathbb{M}_{d}=\mathcal{P}(\Phi)\oplus\mathcal{N}(\Phi)$, where $\mathcal{N}(\Phi)=\{Y:\lim_{n\to\infty}\Phi^{n}(Y)=0\}$ is an orthogonal direct sum with respect to the Hilbert-Schmidt inner product;
  2. (ii)
    $\Phi=\alpha+\beta$, where $\alpha:=\Phi\circ\mathbf{E}_{\mathcal{P}}:\mathbb{M}_{d}\to\mathcal{P}(\Phi)\subseteq\mathbb{M}_{d}\ \text{ and }\ \beta:=\Phi-\alpha=\Phi\circ(\operatorname{\textsf{id}}-\mathbf{E}_{\mathcal{P}}):\mathbb{M}_{d}\to\mathcal{N}(\Phi)\subseteq\mathbb{M}_{d}$;
  3. (iii)
[…]
This shows that the map, $\alpha=\mathrm{Ad}_{U}\circ\mathbf{E}_{\mathcal{P}}=\mathbf{E}_{\mathcal{P}}\circ\mathrm{Ad}_{U}$ is a conditional *-automorphism. Furthermore,
$\left\lVert\Phi^{n}-\alpha^{n}\right\rVert=\left\lVert\beta^{n}\right\rVert\to 0,\penalty 10000\ \penalty 10000\ \penalty 10000\ \text{as }n\to\infty.$
Note that $\mathrm{Ad}_{U}$ is an automorphism of $\mathbb{M}_{d}$ and it leaves $\mathcal{P}(\Phi)$ and $\mathcal{N}(\Phi)$ invariant. Taking $\gamma=\mathrm{Ad}_{U^{*}}$, we see that $\gamma\circ\Phi=\mathbf{E}_{\mathcal{P}}+\mathrm{Ad}_{U^{*}}\circ\beta$. Note that $\mathrm{Ad}_{U^{*}}$ being an automorphism,
$\|(\mathrm{Ad}_{U^{*}}\circ\beta)^{n}\|=\|\beta^{n}\|$
converges to zero. Consequently $(\gamma\circ\Phi)^{n}$ converges to $\mathbf{E}_{\mathcal{P}}$. ∎
Finally we are in a position to state and prove our main result.
Theorem 3.9.
Let $\Phi:\mathbb{M}_{d}\to\mathbb{M}_{d}$ be a unital quantum channel. Then $\{\Phi^{n}\}_{n\in\mathbb{N}}$ is eventually mixed unitary, that is, there exists a $N\geq 1$ such that $\Phi^{n}$ is mixed unitary for all $n\geq N$.
Proof.
We consider the decomposition $\Phi=\alpha+\beta$ and other notation as in Theorem 3.8. Note that $\alpha^{n}=\Phi^{n}\circ\mathbf{E}_{\mathcal{P}}$ and $\mathrm{Ad}_{(U^{*})^{n}}\circ\alpha^{n}=\mathbf{E}_{\mathcal{P}}$ for all $n\geq 1$. Thus by taking $\tau_{n}:=\mathrm{Ad}_{(U^{*})^{n}}\circ\Phi^{n}$ and $X\in\mathcal{P}(\Phi)$, we get
$\tau_{n}(X)$ $=\mathrm{Ad}_{(U^{*})^{n}}\circ\Phi^{n}(X)=\mathrm{Ad}_{(U^{*})^{n}}\circ\Phi^{n}\circ\mathbf{E}_{\mathcal{P}}(X)=\mathrm{Ad}_{(U^{*})^{n}}\circ\alpha^{n}(X)=\mathbf{E}_{\mathcal{P}}(X)=X.$
That is, $\tau_{n}\in\mathcal{F}(\mathcal{P}(\Phi))$ for all $n\geq 1$. Furthermore,
$\left\lVert\beta^{n}\right\rVert$ $=\left\lVert\Phi^{n}-\alpha^{n}\right\rVert=\left\lVert\mathrm{Ad}_{(U^{*})^{n}}\circ(\Phi^{n}-\alpha^{n})\right\rVert=\left\lVert\tau_{n}-\mathbf{E}_{\mathcal{P}}\right\rVert.$
This implies that $\tau_{n}\to\mathbf{E}_{\mathcal{P}}$ as $n\to\infty$. Consider $r_{\mathcal{P}(\Phi)}>0$ as in Theorem 3.4, then we get a $N\geq 1$ such that $\left\lVert\tau^{n}-\mathbf{E}_{\mathcal{P}}\right\rVert\leq r_{\mathcal{P}(\Phi)}$ for all $n\geq N$. Now, by Theorem 3.4, $\tau_{n}$ is mixed unitary for all $n\geq N$. Consequently, $\Phi^{n}=\mathrm{Ad}_{U^{n}}\circ\tau_{n}$ is mixed unitary for all $n\geq N$. ∎
4. Quantum dynamical semigroups of unital quantum channels
Theorem 3.9 establishes that any discrete semigroup of unital quantum channels is eventually mixed unitary. A natural extension is to ask whether an analogous result holds for continuous semigroups of unital quantum channels. To this end, we start by recalling the relevant definitions and results for quantum dynamical semigroups.
Definition 4.1.
A family of linear maps $\{\Phi_{t}\}_{t\geq 0}$ on $\mathbb{M}_{d}$ is said to be a one parameter semigroup if $\Phi_{0}=\operatorname{\textsf{id}};\penalty 10000\ \penalty 10000\ \Phi_{t}\circ\Phi_{s}=\Phi_{s+t}$, for all $t,s\geq 0$, and the map $t\mapsto\Phi_{t}$ is continuous. A one parameter semigroup $\{\Phi_{t}\}$ is said to be a quantum dynamical semigroup if $\Phi_{t}$ is completely positive for every $t\geq 0.$ It is said to be unital $\Phi_{t}(I_{d})=I_{d}$ for every $t$ and a channel if $\Phi_{t}$ is trace-preserving for every $t.$
If $\{\Phi_{t}\}_{t\geq 0}$ is a one parameter semigroup of linear maps on $\mathbb{M}_{d}$, then the map
$\mathcal{L}(X):=\lim_{t\to 0^{+}}\frac{\Phi_{t}(X)-X}{t},\penalty 10000\ \penalty 10000\ X\in\mathbb{M}_{d},$
is a well-defined (bounded) linear map on $\mathbb{M}_{d}$, called the (infinitesimal generator of the semigroup $\{\Phi_{t}\}_{t\geq 0}$. Furthermore,
$\Phi_{t}(X)=e^{t\mathcal{L}}(X)=\sum_{n=0}^{\infty}\frac{(t\mathcal{L})^{n}(X)}{n!},\penalty 10000\ \penalty 10000\ \forall X\in\mathbb{M}_{d},\penalty 10000\ \penalty 10000\ \forall t\geq 0.$
The generators of quantum dynamical semigroups is characterized by the famous GKLS theorem (See [GKS76, Lin76, KRP, CP17]), namely, we have the following theorem:
Theorem 4.2.
[GKS76] Consider a one parameter semigroup of linear maps $\{\Phi_{t}\}_{t\geq 0}$ on $\mathbb{M}_{d}$. Then the following are equivalent:
  1. (i)
    $\{\Phi_{t}\}$ is a quantum dynamical semigroup;
  2. (ii)
    The generator $\mathcal{L}$ of the semigroup $\Phi_{t}$ is of the form
    $\mathcal{L}(X)=G^{*}X+XG+\sum_{j=1}^{n}L_{j}^{*}XL_{j},$
    where $G,L_{j}\in\mathbb{M}_{d},1\leq j\leq n,n\geq 0$.
Further, for a quantum dynamical semigroup $\{\Phi_{t}\},$ with generator $\mathcal{L}$ as above: (a) It is unital iff $\mathcal{L}(I)=0$ iff $G=-iH-\frac{1}{2}\sum_{j=1}^{n}L_{j}^{*}L_{j}$ for some self-adjoint $H$; (b) It is a quantum channel iff $\operatorname{\textsf{tr}}\big(\mathcal{L}(X)\big)=0$ for every $X\in\mathbb{M}_{d}$ iff $G=-iH-\frac{1}{2}\sum_{j=1}^{n}L_{j}L_{j}^{*}.$
In this article we are only considering unital quantum channels. It follows from the previous theorem that a linear map $\mathcal{L}:\mathbb{M}_{d}\to\mathbb{M}_{d}$ is the generator of a one parameter semigroup of unital quantum channels iff it has the form:
$\mathcal{L}(X)=i[H,X]-\frac{1}{2}\sum_{j=1}^{n}(L_{j}^{*}L_{j}X+XL_{j}^{*}L_{j}-2L_{j}^{*}XL_{j}),\penalty 10000\ \penalty 10000\ \forall X\in\mathbb{M}_{d},$ (4.1)
for some $H,L_{j},1\leq j\leq n$ in $\mathbb{M}_{d},\penalty 10000\ n\geq 0,$ satisfying $H=H^{*},\penalty 10000\ \sum_{j=1}^{n}L_{j}L_{j}^{*}=\sum_{j=1}^{n}L_{j}^{*}L_{j}.$
In the case of the continuous dynamical semigroup of mixed unitary channels, the characterization of generators is known in [KuMa87]. They also gave an example of a (continuous) dynamical semigroup of unital quantum channels that is not mixed unitary for some time (see Example 5.10). So, not every dynamical semigroup of unital quantum channels is mixed unitary for all times. Naturally, one can ask what is the behaviour of this semigroup at infinity. More precisely, we are interested in the question: For a given (continuous) dynamical semigroup $\{\Phi_{t}\}_{t\geq 0}$ of unital quantum channels on $\mathbb{M}_{d}$, does there exists a time $t_{0}>0$ such that for all times $t\geq t_{0}$, $\Phi_{t}$ becomes mixed unitary? In this article, we prove that the answer to the above question is affirmative (see Theorem 4.7).
Theorem 4.3.
[KuMa87] Let $\{\Phi_{t}\}_{t\geq 0}$ be a quantum dynamical semigroup of unital channels. Then the following are equivalent:
  1. (i)
    $\Phi_{t}$ is mixed unitary, for all $t\geq 0$;
  2. (ii)
    The generator $\mathcal{L}$ of $\{\Phi_{t}\}$ is given by
    $\mathcal{L}(X)=i[H,X]-\frac{1}{2}\sum_{j=1}^{n}(L_{j}^{2}X+XL_{j}^{2}-2L_{j}XL_{j})+\sum_{k=1}^{\ell}\lambda_{k}(U_{k}^{*}XU_{k}-X),$
    where $H,L_{j}$ are Hermitian, $U_{k}$ are unitary, $\lambda_{k}\geq 0$ with $\sum_{k=1}^{\ell}\lambda_{k}=1$ and $n,\ell\geq 0$;
  3. (iii)
    The infinitesimal generator $\mathcal{L}$ is in the closed cone generated by $\{\mathrm{Ad}_{U}-\operatorname{\textsf{id}}:U\in\mathbb{U}(d)\}$.
We observe that for a generator $\mathcal{L}$ to generate mixed unitary channels at all times it is necessary and sufficient that the coefficients $L_{j}$’s appearing in its GKLS expansion can be taken to be either self-adjoint or scalar multiples of unitary.
In the following, for given $\lambda\in\mathbb{C}$, the Jordan block of size $r\in\mathbb{N}$ with eigenvalue $\lambda$ is defined and denoted as
$J_{r}(\lambda)=\begin{bmatrix}\lambda&1&0&\cdots&0&0\\ 0&\lambda&1&\cdots&0&0\\ \vdots&\vdots&\vdots&\cdots&\vdots&\vdots\\ 0&0&0&\cdots&\lambda&1\\ 0&0&0&\cdots&0&\lambda\end{bmatrix}\in\mathbb{M}_{r}.$
Note that for $r=1$, $J_{1}(\lambda)=[\lambda]\in\mathbb{M}_{1}$. The following Lemma is perhaps known in the literature, but we are providing a proof for completeness.
Lemma 4.4.
Let $\mathcal{L}:\mathbb{M}_{d}\to\mathbb{M}_{d}$ be a linear map and let $\bigoplus_{k=1}^{m}J_{d_{k}}(\lambda_{k})$ be the Jordan decomposition of $\mathcal{L}$ w.r.t. some basis of $\mathbb{M}_{d}$. If $e^{t\mathcal{L}}$ is a quantum dynamical semigroup of unital maps on $\mathbb{M}_{d}$,
  1. (i)
    For every $k$, Re $(\lambda_{k})\leq 0;$
  2. (ii)
    All Jordan blocks corresponding to purely imaginary eigenvalues are one dimensional, that is, if Re $(\lambda_{k})=0$, then $d_{k}=1;$
  3. (iii)
    If Re $(\lambda_{k})<0$, then $\lim_{t\to\infty}e^{tJ_{d_{k}}(\lambda_{k})}=0$.
Proof.
Assume that $\{e^{t\mathcal{L}}\}_{t\geq 0}$ is a quantum dynamical semigroup of unital maps. In particular, it is contractive for every $t$. Since any two norms on $\mathbb{M}_{d}$ are equivalent, it follows that the matrix entries of $e^{t\mathcal{L}}$ are bounded in the Jordan basis of $\mathcal{L}$ as $t$ varies in $0,\infty).$ Fix $k$ and consider the $k$-th Jordan block. If $d_{k}>1$, write $J_{d_{k}}(\lambda_{k})=\lambda I_{d_{k}}+N_{k}$, where $N_{k}$ is the non-zero nilpotent matrix of order $d_{k}$. Then,
$e^{tJ_{d_{k}}(\lambda)}$ $=e^{t\lambda I_{d_{k}}+N_{k}}$
$=e^{t\lambda I_{d_{k}}}e^{tN_{k}}$
$=e^{t\lambda}I_{d_{k}}\big(I_{d_{k}}+tN_{k}+\frac{t^{2}N_{k}^{2}}{2!}+\cdots+\frac{t^{d_{k}-1}N_{k}^{d_{k}-1}}{(d_{k}-1)!}\big).$
Hence
$e^{tJ_{d_{k}}(\lambda_{k})}]_{ij}=\left\{\begin{array}[]{cl}\frac{t^{j-i}}{(j-i)!}\penalty 10000\ e^{t\lambda_{k}}&\qquad i\leq j;\\ 0&\qquad i>j.\end{array}\right$.
Now (i)-(iii) follow easily. ∎
Lemma 4.5.
Let $\mathcal{L}:\mathbb{M}_{d}\to\mathbb{M}_{d}$ be the generator of a quantum dynamical semigroup of unital channels. Then the space $\mathcal{P}:=\operatorname{\textsf{span}}\{X\in\mathbb{M}_{d}:\mathcal{L}(X)=iaX,\quad a\in\mathbb{R}\}$ is a $C^{*}$-sub algebra of $\mathbb{M}_{d}$.
Proof.
To prove $\mathcal{P}$ is a $C^{*}$-subalgebra of $\mathbb{M}_{d}$, it is enough to show that it is closed under multiplication and taking adjoints. Let $X,Y\in\mathbb{M}_{d}$ and $a,b\in\mathbb{R}$ be such that $\mathcal{L}(X)=iaX$ and $\mathcal{L}(Y)=ibY$. Then for all $t\geq 0$,
$e^{t\mathcal{L}}(X)=e^{ita}X\quad\text{and}\quad e^{t\mathcal{L}}(Y)=e^{itb}Y$
This shows that $X,Y$ is in the peripheral space of $e^{t\mathcal{L}}$. Now, as the peripheral space of $e^{t\mathcal{L}}$ lies in the multiplicative domain of $e^{t\mathcal{L}}$, we get that
$e^{t\mathcal{L}}(XY)=e^{t\mathcal{L}}(X)e^{t\mathcal{L}}(Y)=e^{ita+itb}XY,\penalty 10000\ \penalty 10000\ \forall t\geq 0.$
By differentiating both sides at $t=0$, we obtain $\mathcal{L}(XY)=i(a+b)XY$. This shows that it is closed under multiplication. Now, $\mathcal{P}$ is closed under taking adjoints because $\mathcal{L}$ is *-preserving. This completes the proof. ∎
The following is a one parameter semigroup version of Theorem 3.1 of [BhSa23]. For more on the peripheral Poisson boundary of quantum dynamical semigroups one may see [BD25].
Theorem 4.6.
Let $\{\Phi_{t}=e^{t\mathcal{L}}\}_{t\geq 0}$ be a quantum dynamical semigroup of unital channels on $\mathbb{M}_{d}$. Take $\mathcal{P}=\operatorname{\textsf{span}}\{X\in\mathbb{M}_{d}:\mathcal{L}(X)=iaX,\penalty 10000\ \mbox{for some}\penalty 10000\ a\in\mathbb{R}\}$. Then the following hold:
  1. (i)
    $\mathbb{M}_{d}=\mathcal{P}\oplus\mathcal{N}$, where $\mathcal{N}=\{Y:\lim_{t\to\infty}e^{t\mathcal{L}}(Y)=0\}$;
  2. (ii)
    $\Phi_{t}$ leaves $\mathcal{P}$ invariant for every $t$ and restricted to $\mathcal{P}$, $\{\Phi_{t}\}_{t\geq 0}$ form a semigroup of trace-preserving *-automorphisms;
  3. (iii)
    $\Phi_{t}=\alpha_{t}+\beta_{t}$, where $\alpha_{t}:=\Phi_{t}\circ\mathbf{E}_{\mathcal{P}}:\mathbb{M}_{d}\to\mathcal{P}\subseteq\mathbb{M}_{d}$ where $\mathbf{E}_{\mathcal{P}}$ is the unique trace-preserving conditional expectation onto $\mathcal{P}$ and $\beta_{t}:=\Phi_{t}-\alpha_{t}:\mathbb{M}_{d}\to\mathcal{N}\subseteq\mathbb{M}_{d}$ satisfies $\lim_{t\to\infty}\beta_{t}=0$.
Proof.
$(i)$ Using the Jordan decomposition of the generator $\mathcal{L}$ of $\{\Phi_{t}\}$, we write
$\mathbb{M}_{d}=\textsf{ker}(\mathcal{L}-\lambda_{1}\operatorname{\textsf{id}})^{\nu(\lambda_{1})}\oplus\textsf{ker}(\mathcal{L}-\lambda_{2}\operatorname{\textsf{id}})^{\nu(\lambda_{2})}\oplus\cdots\oplus\textsf{ker}(\mathcal{L}-\lambda_{n}\operatorname{\textsf{id}})^{\nu(\lambda_{n})},$
where $\sigma(\mathcal{L})=\{\lambda_{1},\lambda_{2},\cdots,\lambda_{n}\}$ is the set of eigenvalues of $\mathcal{L}$ and $\nu(\lambda_{i})$ is the algebraic multiplicity of the eigenvalue $\lambda_{i}$. From Lemma 4.4, if the real part of $\lambda_{i}=0$, all the Jordan blocks corresponding to the eigenvalue $\lambda_{i}$ are of dimension one and hence the algebraic multiplicity is the same as its geometric multiplicity. Therefore,
$\mathcal{P}=\bigoplus_{\lambda\in\sigma_{pi}(\mathcal{L})}\textsf{ker}(\mathcal{L}-\lambda\operatorname{\textsf{id}})^{\nu(\lambda)},$
where $\sigma_{pi}(\mathcal{L})$ be the set of all purely imaginary eigenvalues of $\mathcal{L}$. Now, if $\lambda_{k}\in\sigma(L)$ be such that real part of $\lambda_{k}\neq 0$, then again by Lemma 4.4, we have $e^{tJ_{k}(\lambda_{k})}\to 0$ as $t\to\infty$. Thus, for $Y\in\textsf{ker}(\mathcal{L}-\lambda_{i}\operatorname{\textsf{id}})^{\nu(\lambda_{i})}$, we have $e^{t\mathcal{L}}(Y)\to 0$ as $t\to\infty$. By letting $\mathcal{N}:=\bigoplus_{\lambda\in\sigma(\mathcal{L})\setminus\sigma_{pi}(\mathcal{L})}\textsf{ker}(\mathcal{L}-\lambda\operatorname{\textsf{id}})^{\nu(\lambda)}$ we get the required decomposition.
[…]
$\operatorname{\textsf{tr}}(X^{*}X)=\operatorname{\textsf{tr}}(e^{t\mathcal{L}}(X^{*}X))=\operatorname{\textsf{tr}}(e^{t\mathcal{L}}(X^{*})e^{t\mathcal{L}}(X))=0.$
[…]
$(iii)$ For every $t\geq 0$, let $\alpha_{t}:=e^{t\mathcal{L}}\circ\mathbf{E}_{\mathcal{P}}$ and $\beta_{t}:=e^{t\mathcal{L}}-\alpha_{t}$. Given $Z\in\mathbb{M}_{d}$, decompose it as $Z=X+Y$ with $X\in\mathcal{P}$ and $Y\in\mathcal{N}$. Then for all $t\geq 0$,
$\beta_{t}(Z)=e^{t\mathcal{L}}(Z)-\alpha_{t}(Z)=e^{t\mathcal{L}}(X)+e^{t\mathcal{L}}(Y)-e^{t\mathcal{L}}\circ\mathbf{E}_{\mathcal{P}}(Z)=e^{t\mathcal{L}}(Y).$
Since $e^{t\mathcal{L}}(Y)\to 0$ as $t\to\infty$, we have $\beta_{t}(\mathbb{M}_{d})\subseteq\mathcal{N}$ for all $t\geq 0$ and $\left\lVert e^{t\mathcal{L}}-\alpha_{t}\right\rVert=\left\lVert\beta_{t}\right\rVert\to 0$ as $t\to\infty$. ∎
We note that part of (i) of the previous theorem does not require $\Phi_{t}$ to be trace-preserving.
Theorem 4.7.
Every quantum dynamical semigroup of unital channels on $\mathbb{M}_{d}$ is eventually mixed unitary.
Proof.
Let $\{\Phi_{t}\}_{t\geq 0}$ be a quantum dynamical semigroup of unital channels on $\mathbb{M}_{d}$. We use the notation and results of Theorem 4.6. Now, from Theorem 3.6, there exists unitaries $\{U_{t}\}_{t\geq 0}\subseteq\mathbb{U}(d)$ such that
[…]
Take $\tau_{t}:=\mathrm{Ad}_{U_{t}^{*}}\circ\Phi_{t}$. Then for every $X\in\mathcal{P}$,
$\tau_{t}(X)$ $=\mathrm{Ad}_{U_{t}^{*}}\circ\Phi_{t}(X)=\mathrm{Ad}_{U_{t}^{*}}\circ\mathrm{Ad}_{U_{t}}(X)=X.$
Note that, $\mathbf{E}_{\mathcal{P}}=\mathrm{Ad}_{U_{t}^{*}}\circ\alpha_{t}$ for all $t\geq 0$, therefore
$\left\lVert\beta_{t}\right\rVert$ $=\left\lVert\Phi_{t}-\alpha_{t}\right\rVert=\left\lVert\mathrm{Ad}_{U_{t}^{*}}\circ(\Phi_{t}-\alpha_{t})\right\rVert=\left\lVert\tau_{t}-\mathbf{E}_{\mathcal{P}}\right\rVert.$
Let $r_{\mathcal{P}}>0$ be as in the Theorem 3.4. Since $\beta_{t}\to 0$ as $t\to\infty$, there exists a $t_{0}\geq 0$ such that $\left\lVert\tau_{t}-\mathbf{E}_{\mathcal{P}}\right\rVert\leq r_{\mathcal{P}}$, for all $t\geq t_{0}$. From Theorem 3.4 (ii), it then follows that $\tau_{t}$ is mixed unitary for all $t\geq t_{0}$. Consequently, $\Phi_{t}=\mathrm{Ad}_{U_{t}}\circ\tau_{t}$ is mixed unitary for all $t\geq t_{0}$. ∎
Definition 4.8.
[HRS20] For any unital trace-preserving linear map $\Phi:\mathbb{M}_{d}\to\mathbb{M}_{d}$, the mixed unitary time or mixed unitary index of $\Phi$ is defined as the least $n\in\mathbb{N}$ such that $\Phi^{k}$ is mixed unitary for all $k\geq n$.
From Theorem 3.9, it follows that the mixed unitary index of a unital quantum channel is finite. One can ask whether the mixed unitary index is independent of $\Phi$. i.e., Is there an integer $n$ such that for every unital quantum channel $\Phi$ on $\mathbb{M}_{d}$ we have $\Phi^{k}$ is mixed unitary, $\forall k\geq n$. The answer is negative. That is, there is no universal power $n$ such that $\Phi^{n}$ mixed unitary for all unital quantum channel $\Phi$ on $\mathbb{M}_{d}$.
Theorem 4.9.
Let $d\geq 3$, then there is no $n$ such that $\Phi^{n}$ is mixed unitary for all unital quantum channels $\Phi:\mathbb{M}_{d}\to\mathbb{M}_{d}$.
Proof.
Suppose that there exists $n\geq 2$ such that $\Phi^{n}$ is mixed unitary for all unital quantum channels $\Phi$ on $\mathbb{M}_{d}$. Let $\mathcal{L}$ be generator of a quantum dynamical semigroup of unital channels, which is not mixed unitary for some $t_{0}>0$. (Note that such $\mathcal{L}$ exists. For example, take $\mathcal{L}$ as in the Lemma 5.9.) Consider $\Phi:=e^{\frac{t_{0}}{n}\mathcal{L}}$. Then $\Phi$ is a unital quantum channel on $\mathbb{M}_{d}$, but $\Phi^{n}=e^{t_{0}\mathcal{L}}$ is not mixed unitary, which gives a contradiction. ∎
5. Behavior of dynamical semigroup of quantum channels near origin
Given any dynamical semigroup of unital quantum channels on $\mathbb{M}_{d}$, we have shown that ( see Theorem 4.7) it is eventually mixed unitary. In this section, we examine the behavior of such semigroups near $t=0$. We do this by first giving a new characterization of generators of semigroups having only mixed unitary channels.
Notations:
$\mathscr{W}(d):$ $=\{\Phi\in\mathscr{H}(d):\Phi\text{ is unital and trace-preserving}\}.$ (5.1)
$\mathscr{W}_{1}(d):$ $=\{\Phi\in\mathscr{W}(d):\left\lVert\Phi\right\rVert=1\}.$ (5.2)
[…]
$\mathscr{D}(d):$ $=\{\Gamma\in\mathscr{W}(d):\langle\operatorname{\textsf{id}},\Gamma\rangle=0,\langle\mathrm{Ad}_{U},\Gamma\rangle\geq 0,\penalty 10000\ \penalty 10000\ \forall U\in\mathbb{U}(d)\}$ (5.4)
In other words, $\mathscr{W}(d)$ is the set of unital, trace-preserving Hermitian maps on $\mathbb{M}_{d}$ and $\mathscr{W}_{1}(d)$ is the collection of unit elements in it. Further, $\mathscr{D}(d)$ is a convenient subset of the dual of the cone generated by mixed unitary channels. Note that the inner product here is defined using Choi matrices as in 2.2. This will be used in the characterization below in Theorem 5.2.
Consider the set $\mathcal{C}(d)^{\circ}\cap\mathscr{W}_{1}(d)$. Then, it is a non-empty compact set. For, since the completely depolarizing channel $\delta_{d}\in\mathscr{W}_{1}(d)$ and $\langle\Phi,\delta_{d}\rangle=\frac{1}{d^{2}}\operatorname{\textsf{tr}}(\mathscr{J}(\Phi))\geq 0$, for all $\Phi\in\mathcal{C}(d)$, we have $\delta_{d}\in\mathcal{C}(d)^{\circ}\cap\mathscr{W}_{1}(d)$. It is compact because it is a closed subset of the compact set $\mathscr{W}_{1}(d)$.
Lemma 5.1.
Suppose $\mathcal{L}\in\mathscr{H}(d)$ with $\langle\Gamma,\mathcal{L}\rangle\geq 0$ for all $\Gamma\in\mathscr{D}(d)$. Then, for every $r>0$ there exists $\eta>0$ such that if $\Gamma\in\mathcal{C}(d)^{\circ}\cap\mathscr{W}_{1}(d)$ satisfies $\langle\Gamma,\operatorname{\textsf{id}}\rangle<\eta$ then $\langle\Gamma,\mathcal{L}\rangle\geq-r$.
Proof.
We will prove by contradiction. Suppose there exists a $r>0$ and a sequence $\{\Gamma_{n}\}$ $\mathcal{C}(d)^{\circ}\cap\mathscr{W}_{1}(d)$ with $\langle\Gamma_{n},\operatorname{\textsf{id}}\rangle<\frac{1}{n}$ such that
$\langle\Gamma_{n},\mathcal{L}\rangle<-r,$ (5.5)
for every $n.$ Since $\mathcal{C}(d)^{\circ}\cap\mathscr{W}_{1}(d)$ is compact, there exist a $\Gamma\in\mathcal{C}(d)^{\circ}\cap\mathscr{W}_{1}(d)$ and a sub sequence $\{\Gamma_{n_{k}}\}$ such that $\lim_{n_{k}}\Gamma_{n_{k}}=\Gamma$ and
$0\leq\langle\Gamma,\operatorname{\textsf{id}}\rangle=\lim_{n_{k}}\langle\Gamma_{n_{k}},\operatorname{\textsf{id}}\rangle<\lim_{n_{k}}\frac{1}{n_{k}}=0.$
This shows that $\Gamma\in\mathscr{D}(d)$. Now, from equation (5.5), we have $\langle\Gamma,\mathcal{L}\rangle\leq-r<0$. This contradicts the assumption on $\mathcal{L}$. ∎
Recall that Kümmerer et al. [KuMa87] characterizes (see Theorem 4.3) the dynamical semigroups that are mixed unitary channels for all $t\geq 0$. Here, we provide another necessary and sufficient condition for a semigroup to be mixed unitary.
Theorem 5.2.
Let $\mathcal{L}\in\mathscr{H}(d)$ be such that $\mathcal{L}(I_{d})=0$ and $\operatorname{\textsf{tr}}(\mathcal{L}(X))=0,\forall X\in\mathbb{M}_{d}$. Then, the following are equivalent:
  1. (i)
    $e^{t\mathcal{L}}$ is mixed unitary for all $t\geq 0$;
  2. (ii)
    The generator $\mathcal{L}$ is in the closed cone generated by $\{\mathrm{Ad}_{U}-\operatorname{\textsf{id}}:U\in\mathbb{U}(d)\}$;
  3. (iii)
    $\langle\Gamma,\mathcal{L}\rangle\geq 0$ for all $\Gamma\in\mathscr{D}(d)$.
Proof.
$(i)\Rightarrow(ii)$ Assume that $e^{t\mathcal{L}}$ is mixed unitary for all $t$. That is, $e^{t\mathcal{L}}=\sum_{j=1}^{n(t)}\lambda_{j}(t)\mathrm{Ad}_{U_{j}(t)}$, for some unitaries $U_{j}(t)$ and scalars $\lambda_{j}(t)\in[0,1]$ with $\sum_{j=1}^{n(t)}\lambda_{j}(t)=1$. Therefore,
$\frac{e^{t\mathcal{L}}-\operatorname{\textsf{id}}}{t}=\sum_{j=1}^{n(t)}\lambda_{j}(t)(\mathrm{Ad}_{U_{j}(t)}-\operatorname{\textsf{id}})\in\text{cone}\{\mathrm{Ad}_{U}-\operatorname{\textsf{id}}:U\in\mathbb{U}(d)\},$
Consequently, $\mathcal{L}=\lim_{t}\frac{e^{t\mathcal{L}}-\operatorname{\textsf{id}}}{t}$ is in the closed cone generated by $\{\mathrm{Ad}_{U}-\operatorname{\textsf{id}}:U\in\mathbb{U}(d)\}$.
$(ii)\Rightarrow(iii)$ Assume that $\mathcal{L}$ is in the closed cone generated by $\{\mathrm{Ad}_{U}-\operatorname{\textsf{id}}:U\in\mathbb{U}(d)\}$. Let $\Gamma\in\mathscr{D}(d)$ be arbitrary. Then, for any unitary $U\in\mathbb{U}(d)$,
$0\leq\langle\Gamma,\mathrm{Ad}_{U}\rangle=\langle\Gamma,\mathrm{Ad}_{U}\rangle-\langle\Gamma,\operatorname{\textsf{id}}\rangle=\langle\Gamma,\mathrm{Ad}_{U}-\operatorname{\textsf{id}}\rangle.$
Therefore, for all $\Phi\in\text{cone}\{\mathrm{Ad}_{U}-\operatorname{\textsf{id}}:U\in\mathbb{U}(d)\}$ we have $\langle\Gamma,\Phi\rangle\geq 0$. Consequently we also get $\langle\Gamma,\Phi\rangle\geq 0$, for all $\Phi$ is in the closed cone generated by $\{\mathrm{Ad}_{U}-\operatorname{\textsf{id}}:U\in\mathbb{U}(d)\}$. In particular, we get $\langle\Gamma,\mathcal{L}\rangle\geq 0$.
$(iii)\Rightarrow(i)$ To prove $e^{t\mathcal{L}}\in\mathscr{MU}(d)$ for all $t\geq 0$, it is enough to show that $e^{\mathcal{L}}\in\mathscr{MU}(d)$. This can be further simplified by showing that $(\operatorname{\textsf{id}}+\frac{\mathcal{L}}{n})\in\mathscr{MU}(d),\penalty 10000\ \penalty 10000\ \forall n\geq n_{0}$, for some fixed $n_{0}$, because $\mathscr{MU}(d)$ is closed under compositions and $e^{\mathcal{L}}=\lim_{n\to\infty}(\operatorname{\textsf{id}}+\frac{\mathcal{L}}{n})^{n}$.
Given $\epsilon>0$, let $\mathcal{L}_{\epsilon}:=\mathcal{L}+\epsilon\delta_{d}$, where $\delta_{d}$ is the completely depolarizing channel. Set $r=\frac{\epsilon}{d^{2}}$. As $\langle\Gamma,\mathcal{L}\rangle\geq 0$ for all $\Gamma\in\mathscr{D}(d)$. From Lemma 5.1, there exists a $\eta>0$ such that $\langle\Gamma,\mathcal{L}\rangle\geq-r$ for all $\Gamma\in\mathcal{C}^{\circ}\cap\mathscr{W}_{1}(d)$ with $\langle\Gamma,\operatorname{\textsf{id}}\rangle<\eta$, where $\mathcal{C}$ is the closed cone generated by $\{\mathrm{Ad}_{U}:U\in\mathbb{U}(d)\}$. Let $\Gamma\in\mathcal{C}^{\circ}\cap\mathscr{W}_{1}(d)$ be arbitrary,
Case (i): If $0\leq\langle\Gamma,\operatorname{\textsf{id}}\rangle<\eta$, then
$\langle\Gamma,\operatorname{\textsf{id}}+\frac{\mathcal{L}_{\epsilon}}{n}\rangle$ $=\langle\Gamma,\operatorname{\textsf{id}}\rangle+\frac{1}{n}\big(\langle\Gamma,\mathcal{L}\rangle+\epsilon\langle\Gamma,\delta_{d}\rangle\big)\geq\frac{1}{n}\big(\langle\Gamma,\mathcal{L}\rangle+\epsilon\langle\Gamma,\delta_{d}\rangle\big)\geq\frac{1}{n}(-\frac{\epsilon}{d^{2}}+\epsilon\frac{1}{d^{2}})=0.$
Case (ii): If $\langle\Gamma,\operatorname{\textsf{id}}\rangle\geq\eta$, then
$\langle\Gamma,\operatorname{\textsf{id}}+\frac{\mathcal{L}_{\epsilon}}{n}\rangle$ $=\langle\Gamma,\operatorname{\textsf{id}}\rangle+\frac{1}{n}\big(\langle\Gamma,\mathcal{L}\rangle+\epsilon\langle\Gamma,\delta_{d}\rangle\big)$
$\geq\eta+\frac{1}{n}\big(\langle\Gamma,\mathcal{L}\rangle+\epsilon\langle\Gamma,\delta_{d}\rangle\big)$
$\geq\eta+\frac{1}{n}\langle\Gamma,\mathcal{L}\rangle$
$\geq\eta-\frac{1}{n}\left\lVert\mathcal{L}\right\rVert_{2}.$
Here $\left\lVert\mathcal{L}\right\rVert_{2}^{2}=\langle\mathcal{L},\mathcal{L}\rangle$. Take $n_{0}\geq 1$ such that $n_{0}\eta\geq\left\lVert\mathcal{L}\right\rVert_{2}$, then
$\langle\Gamma,\operatorname{\textsf{id}}+\frac{\mathcal{L}_{\epsilon}}{n}\rangle=\eta-\frac{1}{n}\left\lVert\mathcal{L}\right\rVert_{2}\geq\eta-\frac{1}{n_{0}}\left\lVert\mathcal{L}\right\rVert_{2}\geq 0,\penalty 10000\ \penalty 10000\ \penalty 10000\ \forall n\geq n_{0}.$ (5.6)
Therefore, for every $n\geq n_{0}$ and $\Gamma\in\mathcal{C}^{\circ}\cap\mathscr{W}_{1}(d)$,
$\langle\Gamma,\operatorname{\textsf{id}}+\frac{\mathcal{L}}{n}\rangle=\lim_{\epsilon\to 0}\langle\Gamma,\operatorname{\textsf{id}}+\frac{\mathcal{L}_{\epsilon}}{n}\rangle\geq 0.$
In particular, $\langle\Gamma,\operatorname{\textsf{id}}+\frac{\mathcal{L}}{n}\rangle\geq 0$ for all $\Gamma\in\mathscr{MU}(d)^{\circ}\cap\mathscr{W}(d)$ and $n\geq n_{0}$. Now, from Lemma 2.7 (iii), it follows that $\operatorname{\textsf{id}}+\frac{\mathcal{L}}{n}$ is mixed unitary for all $n\geq n_{0}$. This completes the proof.∎
Remark 5.3.
From the above Theorem 5.2, it follows that if the infinitesimal generator $\mathcal{L}$ has the form $\Psi-\operatorname{\textsf{id}}$, where $\Psi$ is a mixed unitary channel, then the semigroup $e^{t\mathcal{L}}$ is mixed unitary for all $t\geq 0$. Interestingly, there exist semigroups of mixed unitary channels whose infinitesimal generator is not of this form. As an example, consider the linear map $\mathcal{L}:\mathbb{M}_{3}\to\mathbb{M}_{3}$ given by
$\mathcal{L}(X)=\frac{1}{2}(\operatorname{\textsf{tr}}(X)I_{3}-X^{\operatorname{\textsf{T}}})-X,\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \forall X\in\mathbb{M}_{3}.$
Since the adjiont of the map $X\mapsto\frac{1}{2}(\operatorname{\textsf{tr}}(X)I_{3}-X^{\operatorname{\textsf{T}}})$ is itself, there exists Hermitian matrices $A_{j}\in\mathbb{M}_{3}$ such that $\sum_{j}A_{j}^{2}=I_{3}$ and
$\mathcal{L}(X)=\sum_{j=1}^{\ell}A_{j}XA_{j}-X=\sum_{j=1}^{\ell}\big(A_{j}XA_{j}-\frac{1}{2}(A_{j}^{2}X+XA_{j}^{2})\big),\penalty 10000\ \penalty 10000\ \forall X\in\mathbb{M}_{3}$
By Theorem 4.3, the semigroup $e^{t\mathcal{L}}$ is mixed unitary for all $t\geq 0$.
Theorem 5.4.
Let $\{\Phi_{t}\}_{t\geq 0}$ be a quantum dynamical semigroup of unital channels. Then the set $\{t\in[0,\infty):\Phi_{t}\text{ is not mixed unitary}\}$ is open in $\mathbb{R}$. Moreover one of the following happens:
  1. (i)
    $\Phi_{t}$ is mixed unitary for all $t\geq 0$;
  2. (ii)
    There exist $0<t_{0}\leq t_{1}<\infty$ such that $\Phi_{t}$ is not mixed unitary for $t\in(0,t_{0})$ and is mixed unitary for $t\in[t_{1},\infty)$.
Proof.
Take $N=\{t\in[0,\infty):\Phi_{t}\text{ is not mixed unitary}\}$. Since the set of mixed unitary channels is closed it is obvious that $N$ is open. Let $\mathcal{L}$ be the the generator of $\{\Phi_{t}\}$ so that $\Phi_{t}=e^{t\mathcal{L}}.$ Suppose that $\Phi_{s}$ is not mixed unitary for some $s>0$, from Theorem 5.2, there exists a $\Gamma\in\mathscr{D}(d)$ such that $\langle\Gamma,\mathcal{L}\rangle<0$. Note that if two matrices are Hermitian, then the trace of their product is a real number. This allows us to define a real valued function $g$ on $0,\infty)$ by $g(t):=\langle\Gamma,e^{t\mathcal{L}}\rangle,\penalty 10000\ \penalty 10000\ t\in[0,\infty).$ Then $g$ is a differentiable function with $g(0)=0$. Furthermore,
$g^{\prime}(0)=\lim_{t\to 0}\frac{\langle\Gamma,e^{t\mathcal{L}}\rangle-\langle\Gamma,\operatorname{\textsf{id}}\rangle}{t}=\langle\Gamma,\mathcal{L}\rangle<0.$
Therefore, there exists a $s_{0}>0$ such that $g(t)<0$ for all $t\in(0,s_{0})$. i.e., $\langle\Gamma,\Phi_{t}\rangle<0$ for all $0<t<s_{0}$. Now, by Lemma 2.7 we get that $\Phi_{t}$ is not mixed unitary for all $t\in(0,s_{0})$.
Let $t_{0}:=\max\{t:\Phi_{s}\mbox{ is not mixed unitary },\forall s\in(0,t)\}$. We claim that $\Phi_{t_{0}}$ is a mixed unitary. In deed this clear from openness of $N$, as otherwise $t_{0}\in N$ and this would mean $t_{0},t_{0}+\epsilon)\subseteq N$ for some $\epsilon>0.$ Now, define
$t_{1}:=\min\{t:\Phi_{s}\mbox{ is mixed unitary },\forall s\in[t,\infty)\}.$
By Theorem 4.7, $t_{1}$ is finite. Clearly $t_{0}\leq t_{1}$. ∎
We wish to remark that in the previous theorem we do not know whether $t_{0}<t_{1}$ is feasible or not. Considering the general features of numerous situations of eventual positivity (See for example [Ja17]) we would guess that this could be possible for some examples with the set of points in $t_{0},t_{1}]$ for which $\Phi_{t}$ is a mixed unitary channel forming a discrete set. We are not able to confirm this in any of the following examples.
We now give an example of a dynamical semigroup of unital quantum channels, which is not mixed unitary for some $t>0$. For any $B\in\mathbb{M}_{d}$, let $\mu(B)$ denote the set of all singular values of $B$.
Lemma 5.5.
[MeWo09, Corollory 15] Given non-negative numbers $\mu_{1}\geq\mu_{2}\geq\cdots\geq\mu_{d}$ we have
$\min_{A}\{\operatorname{\textsf{tr}}(A\overline{A}):\mu(A)=\{\mu_{j}\}_{j=1}^{d}\}$ $=\begin{cases*}-2\Big(\sum_{j=1}^{\frac{d}{2}}\mu_{2j}\mu_{2j-1}\Big)&if $d$ is even\\ -2\Big(\sum_{j=1}^{\frac{d-1}{2}}\mu_{2j}\mu_{2j-1}\Big)+\mu_{d}^{2}&if $d$ is odd\end{cases*}$ (5.7)
Lemma 5.6.
Let $F=\sum_{i,j=1}^{d}E_{ij}\otimes E_{ji}\in\mathbb{M}_{d}\otimes\mathbb{M}_{d}$. For any $B\in\mathbb{M}_{d}$, let $W=(I_{d}\otimes B)F(I_{d}\otimes B)^{*}$. Then
$\min_{A}\{\operatorname{\textsf{tr}}(A\overline{A}):\mu(A)=\mu(B)\}$ $\leq\min\{\operatorname{\textsf{tr}}(W|{(I_{d}\otimes U)e}\rangle\langle{(I_{d}\otimes U)e}|):U\in\mathbb{U}(d)\},$ (5.8)
where $e=\sum_{j=1}^{d}e_{j}\otimes e_{j}$.
Proof.
For every unitary $U\in\mathbb{U}(d)$ we have
$\operatorname{\textsf{tr}}(W|{(I_{d}\otimes U)e}\rangle\langle{(I_{d}\otimes U)e}|)$ $=\langle(I_{d}\otimes U)e,W((I_{d}\otimes U)e)\rangle=\langle\sum_{j}e_{j}\otimes Ue_{j},W(\sum_{k}e_{k}\otimes Ue_{k})\rangle$
$=\sum_{j,k}\langle e_{j}\otimes B^{*}Ue_{j},F(e_{k}\otimes B^{*}Ue_{k})\rangle=\sum_{j,k}\langle e_{j}\otimes B^{*}Ue_{j},B^{*}Ue_{k}\otimes e_{k}\rangle$
$=\sum_{j,k}\langle e_{j},B^{*}Ue_{k}\rangle\langle B^{*}Ue_{j},e_{k}\rangle=\operatorname{\textsf{tr}}(B^{*}U\overline{B^{*}U}).$
Note that if $A=B^{*}U$, then $A$ has the same singular values as $B$. Therefore, we have $\min_{A}\{\operatorname{\textsf{tr}}(A\overline{A}):\mu(A)=\mu(B)\}\leq\operatorname{\textsf{tr}}(W|{(I_{d}\otimes U)e}\rangle\langle{(I_{d}\otimes U)e}|)$ for all $U\in\mathbb{U}(d)$. Consequently,
$\min_{A}\{\operatorname{\textsf{tr}}(A\overline{A}):\mu(A)=\mu(B)\}\leq\min\{\operatorname{\textsf{tr}}(W|{(I_{d}\otimes U)e}\rangle\langle{(I_{d}\otimes U)e}|):U\in\mathbb{U}(d)\}.$
Lemma 5.7.
Let $B\in\mathbb{U}(3)$ be any unitary matrix such that $\operatorname{\textsf{tr}}(B^{*}B^{\operatorname{\textsf{T}}})=-1$, define the linear maps $\mathcal{L},\Gamma$ on $\mathbb{M}_{3}$ by
$\mathcal{L}(X):$ $=\frac{1}{2}\big(\operatorname{\textsf{tr}}(X)I_{3}-BX^{\operatorname{\textsf{T}}}B^{*}-2X\big),$
$\Gamma(X):$ $=\frac{1}{2}\Big(BX^{\operatorname{\textsf{T}}}B^{*}+\frac{1}{3}\operatorname{\textsf{tr}}(X)I_{3}\Big),\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \forall X\in\mathbb{M}_{3}.$
Then the following holds:
  1. (i)
    $\mathcal{L}$ is Hermitian preserving map with $\mathcal{L}(I_{3})=0$ and $\operatorname{\textsf{tr}}(\mathcal{L}(X))=0$.
  2. (ii)
    $\Gamma\in\mathscr{D}(3)$.
  3. (iii)
    $\langle\Gamma,\mathcal{L}\rangle=-\frac{1}{9}$.
Proof.
$(i)$ Follows from the definition of $\mathcal{L}$.
$(ii)$ Clearly, $\Gamma$ is unital and trace-preserving map. Take $W=(I_{d}\otimes B)F(I_{d}\otimes B)^{*}$, where $F=\sum_{i,j=1}^{3}E_{ij}\otimes E_{ji}$. Then $\mathscr{J}(\Gamma)=\frac{1}{6}(W+\frac{I_{3}\otimes I_{3}}{3})$, and
$\langle\mathrm{Ad}_{I_{3}},\Gamma\rangle=\operatorname{\textsf{tr}}(\mathscr{J}(\Gamma)|{\Omega_{3}}\rangle\langle{\Omega_{3}}|)$ $=\frac{1}{6}\operatorname{\textsf{tr}}\big((W+\frac{I_{3}\otimes I_{3}}{3})|{\Omega_{3}}\rangle\langle{\Omega_{3}}|\big)$
$=\frac{1}{6}\big(\operatorname{\textsf{tr}}(W|{\Omega_{3}}\rangle\langle{\Omega_{3}}|)+\frac{1}{3}\big)$
$=\frac{1}{6}\big(\frac{1}{3}\operatorname{\textsf{tr}}(B^{*}B^{\operatorname{\textsf{T}}})+\frac{1}{3}\big)=0.$ $(\because\operatorname{\textsf{tr}}(B^{*}B^{\operatorname{\textsf{T}}})=-1).$
Now, for any unitary $U\in\mathbb{U}(3)$, by using the Lemma 5.6 and 5.5, we have
$\langle\mathrm{Ad}_{U},\Gamma\rangle=\operatorname{\textsf{tr}}(\mathscr{J}(\mathrm{Ad}_{U})\mathscr{J}(\Gamma))$ $=\operatorname{\textsf{tr}}\Big(\mathscr{J}(\mathrm{Ad}_{U})\frac{1}{6}\big(W+\frac{1}{3}I_{3}\otimes I_{3}\big)\Big)$
$=\frac{1}{6}\operatorname{\textsf{tr}}(\mathscr{J}(\mathrm{Ad}_{U})W)+\frac{1}{18}\operatorname{\textsf{tr}}(\mathscr{J}(\mathrm{Ad}_{U}))$
$=\frac{1}{18}\operatorname{\textsf{tr}}(W|{(I_{3}\otimes U^{*})e}\rangle\langle{(I_{3}\otimes U^{*})e}|)+\frac{1}{18}$
$\geq\frac{1}{18}\min\{\operatorname{\textsf{tr}}(W|{(I_{3}\otimes V)e}\rangle\langle{(I_{3}\otimes V)e}|):V\in\mathbb{U}(3)\}+\frac{1}{18}$
$\geq\frac{1}{18}\min\{\operatorname{\textsf{tr}}(A\overline{A}):\mu(A)=\mu(B)\}+\frac{1}{18}$
$=\frac{1}{18}\min\{\operatorname{\textsf{tr}}(A\overline{A}):\mu(A)=(1,1,1)\}+\frac{1}{18}$ $(\because\mu(B)=(1,1,1))$
$=\frac{1}{18}(-2+1)+\frac{1}{18}=0.$
This shows that $\Gamma\in\mathscr{D}(3)$.
$(iii)$ Let $W=(I_{d}\otimes B)F(I_{d}\otimes B)^{*}$, where $F=\sum_{i,j=1}^{3}E_{ij}\otimes E_{ji}$. Let $\eta:\mathbb{M}_{3}\to\mathbb{M}_{3}$ be a linear map given by
$\eta(X):=\frac{1}{2}(\operatorname{\textsf{tr}}(X)I_{3}-X^{\operatorname{\textsf{T}}}),\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \forall X\in\mathbb{M}_{3}.$
Then it is easy to see that $\mathcal{L}=\mathrm{Ad}_{B^{*}}\circ\eta-\operatorname{\textsf{id}}$ and
$\mathscr{J}(\mathrm{Ad}_{B^{*}}\circ\eta)$ $=(I_{3}\otimes B)\mathscr{J}(\eta)(I_{3}\otimes B^{*})=\frac{1}{6}(I_{3}\otimes B)(I_{3}\otimes I_{3}-F)(I_{3}\otimes B^{*})=\frac{1}{6}(I_{3}\otimes I_{3}-W),$
$\mathscr{J}(\Gamma)$ $=\frac{1}{6}(\frac{1}{3}I_{3}\otimes I_{3}+W).$
Therefore,
$\langle\Gamma,\mathcal{L}\rangle=\langle\Gamma,\mathrm{Ad}_{B^{*}}\circ\eta\rangle-\langle\Gamma,\mathrm{Ad}_{I_{3}}\rangle$ $=\operatorname{\textsf{tr}}(\mathscr{J}(\Gamma)\mathscr{J}(\mathrm{Ad}_{B^{*}}\circ\eta))-0$
$=\frac{1}{36}\operatorname{\textsf{tr}}\big((W+\frac{1}{3}I_{3}\otimes I_{3})(I_{3}\otimes I_{3}-W)\big)$
$=\frac{1}{36}\operatorname{\textsf{tr}}\big(W-W^{2}+\frac{1}{3}(I_{3}\otimes I_{3}-W)\big)$
$=\frac{1}{36}\operatorname{\textsf{tr}}(W-I_{3}\otimes I_{3}+\frac{I_{3}\otimes I_{3}}{3}-\frac{W}{3})$ $(\because W^{2}=I_{3}\otimes I_{3})$
$=\frac{1}{36}\frac{2}{3}\operatorname{\textsf{tr}}(W-I_{3}\otimes I_{3})$
$=\frac{1}{54}\big(\operatorname{\textsf{tr}}(W)-\operatorname{\textsf{tr}}(I_{3}\otimes I_{3})\big)$
$=\frac{1}{54}(3-9)=\frac{-6}{54}=\frac{-1}{9}.$
Remark 5.8.
Note that there are unitaries $B\in\mathbb{M}_{3}$ such that $\operatorname{\textsf{tr}}(B^{*}B^{\operatorname{\textsf{T}}})=-1$. For example,
$\frac{1}{2}\begin{bmatrix}0&1-i&-1-i\\ -1+i&-i&1\\ 1+i&1&i\end{bmatrix},\qquad\begin{bmatrix}1&0&0\\ 0&0&-1\\ 0&1&0\end{bmatrix},\qquad\begin{bmatrix}0&-1&0\\ 1&0&0\\ 0&0&1\end{bmatrix}$.
Example 5.9.
Let $\mathcal{L}$ be a linear map on $\mathbb{M}_{3}$ given by
$\mathcal{L}(X)=\frac{1}{2}B(\operatorname{\textsf{tr}}(X)I_{3}-X^{\operatorname{\textsf{T}}})B^{*}-X,\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \forall X\in\mathbb{M}_{3},$
where $B\in\mathbb{M}_{3}$ be any unitary matrix such that $\operatorname{\textsf{tr}}(B^{*}B^{\operatorname{\textsf{T}}})=-1$. Now, consider the semigroup whose infinitesimal generator is $\mathcal{L}$. From Theorem 4.2, it follows that $\{e^{t\mathcal{L}}\}_{t\geq 0}$ is a dynamical semigroup of unital quantum channels. Take $\Gamma=\frac{1}{2}(\mathrm{Ad}_{B^{*}}\circ\operatorname{\textsf{T}}+\delta_{3})$, then from Lemma 5.7, $\Gamma\in\mathscr{D}(3)$ and $\langle\Gamma,\mathcal{L}\rangle=-\frac{1}{9}<0$. Therefore, by Theorem 5.2, the semigroup $e^{t\mathcal{L}}$ is not mixed unitary for some $t>0$.
Example 5.10.
[KuMa87] Let $A\in\mathbb{M}_{d}$ be such that either $A$ is non-normal, or its spectrum is neither contained in a circle nor a straight line. Take the linear map $\mathcal{L}$ on $\mathbb{M}_{d}$
$\mathcal{L}(X)=A^{*}XA-\frac{1}{2}(A^{*}AX+XA^{*}A),\penalty 10000\ \penalty 10000\ \penalty 10000\ \forall X\in\mathbb{M}_{d}.$
Then the semigroup of unital quantum channel $e^{t\mathcal{L}}$ is not mixed unitary for some $t>0$.
We have shown in the above Theorem 5.4, that there exists $t_{0}$ such that $e^{t\mathcal{L}}$ is not mixed unitary for any $t\in(0,t_{0})$. We now give an example of a semigroup for which we determine $t_{0}$ such that $e^{t\mathcal{L}}$ is not mixed unitary for any $t\in(0,t_{0})$.
Example 5.11.
Let $\mathcal{L}$ be the linear map on $\mathbb{M}_{3}$ as given in the Lemma 5.7. That is,
$\mathcal{L}(X)=\frac{1}{2}B(\operatorname{\textsf{tr}}(X)I_{3}-X^{\operatorname{\textsf{T}}})B^{*}-X,\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \forall X\in\mathbb{M}_{3},$
where $B=\begin{bmatrix}1&0&0\\ 0&0&-1\\ 0&1&0\end{bmatrix}$. Consider the semigroup $e^{t\mathcal{L}}$. Then, by Theorem 4.2 $e^{t\mathcal{L}}$ is a dynamical semigroup of unital quantum channels. Let $t_{0}\approx 1.4034$ be the positive root of the function $e^{t}-e^{-\frac{t}{2}}-3\sinh(\frac{t}{2})-2\sin(\frac{t}{2})$.
Claim: $e^{t\mathcal{L}}$ is not mixed unitary, for all $0<t<t_{0}$.
Proof of the claim : For $1\leq j\leq 3$, let $\mathcal{L}_{j}$ be the linear maps on $\mathbb{M}_{3}$ given by
$\mathcal{L}_{1}$ $:=-\frac{1}{2}\mathrm{Ad}_{B^{*}}\circ\operatorname{\textsf{T}},\qquad\mathcal{L}_{2}:=\frac{3}{2}\delta_{3},\qquad\mathcal{L}_{3}:=-\operatorname{\textsf{id}}.$
Then $\mathcal{L}_{j}$’s commute with each other, and
$e^{t\mathcal{L}_{3}}$ $=\lim_{n}(\operatorname{\textsf{id}}+\sum_{k=1}^{n}\frac{(-t)^{k}}{k!}\operatorname{\textsf{id}})=\lim_{n}(1+\sum_{k=1}^{n}\frac{(-t)^{k}}{k!})\operatorname{\textsf{id}}=e^{-t}\operatorname{\textsf{id}},$
$e^{t\mathcal{L}_{2}}$ $=\lim_{n}(\operatorname{\textsf{id}}+\sum_{k=1}^{n}\frac{(3t)^{k}}{2^{k}k!}\delta_{3}^{k})=\lim_{n}(\operatorname{\textsf{id}}+\sum_{k=1}^{n}\frac{(3t)^{k}}{2^{k}k!}\delta_{3})=\operatorname{\textsf{id}}+(e^{\frac{3}{2}t}-1)\delta_{3}.$
To compute $e^{t\mathcal{L}_{1}}$, first note that $\mathrm{Ad}_{B^{\operatorname{\textsf{T}}}}\circ\operatorname{\textsf{T}}=\mathrm{Ad}_{B^{*}}\circ\operatorname{\textsf{T}}=\operatorname{\textsf{T}}\circ\mathrm{Ad}_{B^{\operatorname{\textsf{T}}}}$, $B^{2}=(B^{*})^{2}=(B^{\operatorname{\textsf{T}}})^{2}$ and $B=(B^{*})^{3}=(B^{\operatorname{\textsf{T}}})^{3}$,
[…]
Therefore,
$e^{t\mathcal{L}_{1}}$ $=\lim_{n}(\operatorname{\textsf{id}}+\sum_{k=1}^{n}\frac{(-t)^{k}}{2^{k}k!}(\mathrm{Ad}_{B}\circ\operatorname{\textsf{T}})^{k})$
$=\operatorname{\textsf{id}}+\sum_{k=0\text{mode}4}\frac{(-t)^{k}}{2^{k}k!}\operatorname{\textsf{id}}+\sum_{k=1\text{mode}4}\frac{(-t)^{k}}{2^{k}k!}\mathrm{Ad}_{B^{\operatorname{\textsf{T}}}}\circ\operatorname{\textsf{T}}+\sum_{k=2\text{mode}4}\frac{(-t)^{k}}{2^{k}k!}\mathrm{Ad}_{(B^{\operatorname{\textsf{T}}})^{2}}+\sum_{k=3\text{mode}4}\frac{(-t)^{k}}{2^{k}k!}\mathrm{Ad}_{B}\circ\operatorname{\textsf{T}}$
$=a_{t}\operatorname{\textsf{id}}+b_{t}\mathrm{Ad}_{B^{\operatorname{\textsf{T}}}}\circ\operatorname{\textsf{T}}+c_{t}\mathrm{Ad}_{(B^{\operatorname{\textsf{T}}})^{2}}+d_{t}\mathrm{Ad}_{B}\circ\operatorname{\textsf{T}},$
where
$a_{t}=\sum_{n=0}^{\infty}\frac{t^{4n}}{2^{4n}(4n)!},\quad b_{t}=\sum_{n=0}^{\infty}\frac{(-t)^{4n+1}}{2^{4n+1}(4n+1)!},\quad c_{t}=\sum_{n=0}^{\infty}\frac{t^{4n+2}}{2^{4n+2}(4n+2)!},\quad d_{t}=\sum_{n=0}^{\infty}\frac{(-t)^{4n+3}}{2^{4n+3}(4n+3)!}.$
Furthermore, we have
$\langle\mathrm{Ad}_{(B^{\operatorname{\textsf{T}}})^{2}},\Gamma\rangle$ $=\frac{1}{2}\langle\operatorname{\textsf{id}},\mathrm{Ad}_{B^{2}}\circ(\mathrm{Ad}_{B^{*}}\circ\operatorname{\textsf{T}}+\delta_{3})\rangle=\frac{1}{2}\langle\operatorname{\textsf{id}},\mathrm{Ad}_{B}\circ\operatorname{\textsf{T}}+\delta_{3})\rangle=\frac{1}{2}(\langle\operatorname{\textsf{id}},\mathrm{Ad}_{B}\circ\operatorname{\textsf{T}}\rangle+\langle\operatorname{\textsf{id}},\delta_{3}\rangle)$
$=\frac{1}{2}\big(\operatorname{\textsf{tr}}(\mathscr{J}(\operatorname{\textsf{id}})\mathscr{J}(\mathrm{Ad}_{B}\circ\operatorname{\textsf{T}}))+\operatorname{\textsf{tr}}(\mathscr{J}(\operatorname{\textsf{id}})\mathscr{J}(\delta_{3}))=\frac{1}{6}\big(\langle\Omega_{d},(I_{3}\otimes B^{*})F(I_{3}\otimes B)\Omega_{d}\rangle+\frac{1}{3}\langle\Omega_{d},\Omega_{d}\rangle\big)$
$=\frac{1}{6}\big(\frac{\operatorname{\textsf{tr}}(\overline{B}B)}{3}+\frac{1}{3}\big)=\frac{1}{6}\big(\frac{-1}{3}+\frac{1}{3}\big)=0,$
$\langle\mathrm{Ad}_{B^{\operatorname{\textsf{T}}}}\circ\operatorname{\textsf{T}},\Gamma\rangle=\langle\operatorname{\textsf{T}},\mathrm{Ad}_{B}\circ\Gamma\rangle$ $=\frac{1}{2}\langle\operatorname{\textsf{T}},\mathrm{Ad}_{B}\circ(\mathrm{Ad}_{B^{*}}\circ\operatorname{\textsf{T}}+\delta_{3})\rangle$
$=\frac{1}{2}\big(\langle\operatorname{\textsf{id}},\operatorname{\textsf{id}}+\delta_{3}\rangle\big)=\frac{1}{2}\big(\left\lvert\langle\Omega_{3},\Omega_{3}\rangle\right\rvert^{2}+\frac{1}{9}\langle\Omega_{3},\Omega_{3}\rangle\big)=\frac{1}{2}(1+\frac{1}{9})=\frac{10}{18},$
and
$\langle\mathrm{Ad}_{B}\circ\operatorname{\textsf{T}},\Gamma\rangle=$ $\langle\operatorname{\textsf{T}},\mathrm{Ad}_{B^{*}}\circ\Gamma\rangle=\frac{1}{2}\langle\operatorname{\textsf{T}},\mathrm{Ad}_{B^{*}}\circ(\mathrm{Ad}_{B^{*}}\circ\operatorname{\textsf{T}}+\delta_{3})\rangle$
$=\frac{1}{2}\big(\langle\operatorname{\textsf{T}},\mathrm{Ad}_{(B^{2})^{*}}\circ\operatorname{\textsf{T}}+\delta_{3}\rangle\big)=\frac{1}{2}\big(\langle\operatorname{\textsf{id}},\mathrm{Ad}_{(B^{2})^{*}}\rangle+\langle\operatorname{\textsf{id}},\delta_{3}\rangle\big)$
$=\frac{1}{2}\big(\left\lvert\langle(I_{3}\otimes B^{2})\Omega_{3},\Omega_{3}\rangle\right\rvert^{2}+\frac{1}{9}\langle\Omega_{3},\Omega_{3}\rangle\big)=\frac{1}{2}\big(\frac{1}{9}\left\lvert\operatorname{\textsf{tr}}(\overline{B}^{2})\right\rvert^{2}+\frac{1}{9}\big)=\frac{1}{9}$
Therefore, for all $t\in(0,t_{0})$ we get,
$\langle e^{t\mathcal{L}},\Gamma\rangle$ $=(1-e^{-\frac{3}{2}t})\langle\delta_{3},\Gamma\rangle+e^{-t}\big\{a_{t}\langle\operatorname{\textsf{id}},\Gamma\rangle+b_{t}\langle\mathrm{Ad}_{B^{\operatorname{\textsf{T}}}}\circ\operatorname{\textsf{T}},\Gamma\rangle+c_{t}\langle\mathrm{Ad}_{(B^{\operatorname{\textsf{T}}})^{2}},\Gamma\rangle+d_{t}\langle\mathrm{Ad}_{B}\circ\operatorname{\textsf{T}},\Gamma\rangle\big\}$
$=(1-e^{-\frac{3}{2}t})\frac{1}{9}+e^{-t}(\frac{10}{18}b_{t}+\frac{1}{9}d_{t})$
$=\frac{e^{-t}}{9}\big\{e^{t}-e^{-\frac{t}{2}}+5b_{t}+d_{t}\big\}$
$=\frac{e^{-t}}{9}\big\{e^{t}-e^{-\frac{t}{2}}+4b_{t}-\sinh(\frac{t}{2})\big\}\qquad(\because b_{t}+d_{t}=\sinh(\frac{-t}{2}))$
$=\frac{e^{-t}}{9}\big\{e^{t}-e^{-\frac{t}{2}}-\frac{4}{2}(\sinh(\frac{t}{2})-\sin(\frac{t}{2}))-\sinh(\frac{t}{2})\big\}\qquad(\because b_{t}=-\sinh(\frac{t}{2})-\sin(\frac{t}{2}))$
$=\frac{e^{-t}}{9}\big\{e^{t}-e^{-\frac{t}{2}}-3\sinh(\frac{t}{2})-2\sin(\frac{t}{2})\big\}<0$
Now, from Lemma 2.7, it follows that $e^{t\mathcal{L}}$ is not mixed unitary for all $0<t<t_{0}$.
6. Semigroups of mixed Weyl-unitary channels
A mixed unitary channel is a convex combination of maps of the form $\mathrm{Ad}_{U}$, where $U$ can be arbitrary unitary. Let $\mathscr{S}$ be a subset of the group of all $d\times d$ unitaries $\mathbb{U}(d)$. One can ask, What are all the channels that can be written as convex combinations of maps $\mathrm{Ad}_{U}$, where $U\in\mathscr{S}$? In this context we recall the notion of Weyl unitaries.
Let $\mathbb{Z}_{d}=\{0,1,\ldots,d-1\}$ be the cyclic group of order $d$, and let $\xi=e^{\frac{2\pi i}{d}}$ be the primitive $d$-th root of unity. We consider the Hilbert space $\mathbb{C}^{d}$ with orthonormal basis indexed by $\mathbb{Z}_{d}$: $\{e_{j}:0\leq j\leq(d-1)\}$. For every $a,b\in\mathbb{Z}_{d}$, define the unitary operators, $U,V$, and $W_{a,b}$ on $\mathbb{C}^{d}$ as follows:
$U(e_{j})$ $=e_{j+1};$
$V(e_{j})$ $=\xi^{j}e_{j};$
$W_{a,b}(e_{j})$ $=U^{a}V^{b}(e_{j}),\qquad\penalty 10000\ \forall j,a,b\in\mathbb{Z}_{d}.$
The operators $W_{a,b}$ are known as Weyl unitaries. The usefulness of Weyl unitaries in quantum information theory is well-known (See [Wat18]). They have the following key properties:
  1. (i)
    (Adjoint) $W_{a,b}^{*}=\xi^{ab}W_{-a,-b}$, for all $a,b\in\mathbb{Z}_{d}$;
  2. (ii)
    (Commutation relations) $W_{a,b}W_{a^{\prime},b^{\prime}}=\xi^{ba^{\prime}-ab^{\prime}}W_{a^{\prime},b^{\prime}}W_{a,b}$, for all $a,b,a^{\prime},b^{\prime}\in\mathbb{Z}_{d}$;
  3. (iii)
    (Orthogonality) $\langle W_{a,b},W_{a^{\prime},b^{\prime}}\rangle=\operatorname{\textsf{tr}}(W_{a,b}^{*}W_{a^{\prime},b^{\prime}})=\begin{cases*}d&\qquad\text{if} (a,b)=(a',b')\\ 0&\qquad\text{otherwise}\end{cases*}$.
The orthogonality condition tells us that, the set $\{\frac{1}{\sqrt{d}}W_{a,b}:a,b\in\mathbb{Z}_{d}\}$ is an orthonormal basis for $\mathbb{M}_{d}$ w.r.t. the Hilbert-Schmidt inner product. For any $a,b,a^{\prime},b^{\prime}\in\mathbb{Z}_{d}$, define $\phi_{(a,b),(a^{\prime},b^{\prime})}$ on $\mathbb{M}_{d}$ as
$\phi_{(a,b),(a^{\prime},b^{\prime})}(X)=W_{a,b}XW_{a^{\prime},b^{\prime}}^{*}\qquad\penalty 10000\ \forall X\in\mathbb{M}_{d}.$
Observe that the family $\{\phi_{(a,b),(a^{\prime},b^{\prime})}:a,b,a^{\prime},b^{\prime}\in\mathbb{Z}_{d}\}$ forms a basis for $\mathscr{B}({\mathbb{M}_{d}})$. Therefore, any linear map $\Phi$ on $\mathbb{M}_{d}$ is written as $\Phi=\sum_{a,b,a^{\prime},b^{\prime}\in\mathbb{Z}_{d}}\lambda_{(a,b),(a^{\prime},b^{\prime})}\phi_{(a,b),(a^{\prime},b^{\prime})}$. We let $\lambda_{(a,b),(a,b)}=\lambda_{a,b}$ and $\phi_{(a,b),(a,b)}=\phi_{a,b}$.
Definition 6.1.
A linear map $\Phi:\mathbb{M}_{d}\to\mathbb{M}_{d}$ is said to be
  1. (i)
    mixed Weyl-unitary if there exist positive scalars $\lambda_{1},\lambda_{2},\ldots,\lambda_{l}$ add up to $1$ and Weyl-unitaries $U_{1},U_{2},\ldots,U_{l}$ in $\mathbb{M}_{d}$ such that $\Phi(X)=\sum_{i=1}^{l}\lambda_{i}U_{i}XU_{i}^{*}$ for all $X\in\mathbb{M}_{d}$.
  2. (ii)
    Weyl-covariant if $\Phi(W_{a,b}XW_{a,b}^{*})=W_{a,b}\Phi(X)W_{a,b}^{*}$ for all $X\in\mathbb{M}_{d}$ and $a,b\in\mathbb{Z}_{d}$.
Proposition 6.2.
[Wat18] Let $\Phi$ be a unital quantum channel on $\mathbb{M}_{d}$. Then the following are equivalent:
  1. (i)
    $\Phi$ is a mixed Weyl-unitary;
  2. (ii)
    $\Phi$ is Weyl-covariant;
  3. (iii)
    $\Phi=\sum_{a,b\in\mathbb{Z}_{d}}\lambda_{a,b}\phi_{a,b}$ with $\lambda_{a,b}\geq 0$ such that $\sum_{a,b\in\mathbb{Z}_{d}}\lambda_{a,b}=1$.
Lemma 6.3.
The following holds:
  1. (i)
    The set of all mixed Weyl-unitaries is a compact, convex set of the real vector space $\mathscr{H}(d)$.
  2. (ii)
    The completely depolarizing channel $\delta_{d}$ is mixed Weyl-unitary.
  3. (iii)
    For a unital channel $\Phi$ on $\mathbb{M}_{d}$ we have: $\Phi$ is mixed Weyl-unitary if and only if $\langle\Gamma,\Phi\rangle\geq 0$ for all unital and trace-preserving $\Gamma\in\mathscr{H}(d)$ with $\langle\Gamma,\mathrm{Ad}_{U}\rangle\geq 0$ for all Weyl-unitaries $U\in\mathbb{U}(d)$.
Proof.
$(i)$ It is easy to see that the set of all Weyl-covariant channels is a closed and convex set of the real vector space $\mathscr{H}(d)$. Now, from Proposition 6.2, the set of mixed Weyl-unitaries is equal to the set of unital Weyl-covariant channels. Thus, the set of mixed Weyl-unitaries is a closed (hence compact) convex set.
$(ii)$ Since $\{\frac{1}{\sqrt{d}}W_{a,b}:a,b\in\mathbb{Z}_{d}\}$ is an orthonormal basis of $\mathbb{M}_{d}$ w.r.t. the Hilbert-Schmidt inner product, we have
$I_{d}\otimes I_{d}=\sum_{a,b\in\mathbb{Z}_{d}}\frac{1}{d}|{w_{a,b}}\rangle\langle{w_{a,b}}|,$ (6.1)
where $w_{a,b}:=\sum_{j\in\mathbb{Z}_{d}}e_{j}\otimes W_{a,b}^{*}e_{j}$. Observe that $\frac{1}{d}|{w_{a,b}}\rangle\langle{w_{a,b}}|=\mathscr{J}(\mathrm{Ad}_{W_{a,b}})$ and $\frac{1}{d^{2}}I_{d}\otimes I_{d}=\mathscr{J}(\delta_{d})$. Therefore, by applying $\mathscr{J}^{-1}$ to the equation (6.1), we get $\delta_{d}=\sum_{a,b\in\mathbb{Z}_{d}}\frac{1}{d^{2}}\mathrm{Ad}_{W_{a,b}}.$
$(ii)$ Proof follows similar to the proof of 2.7 (iii). ∎
Theorem 6.4.
Let $e^{t\mathcal{L}}$ be a dynamical semigroup of unital quantum channels. Then the following are equivalent:
  1. (i)
    $e^{t\mathcal{L}}$ is mixed Weyl-unitary for every $t\geq 0$;
  2. (ii)
    $\mathcal{L}$ is a non-negative linear combination of maps of the form $\mathrm{Ad}_{W}-\operatorname{\textsf{id}}$, where $W$ is a Weyl unitary;
  3. (iii)
    $\mathcal{L}$ is a Weyl-covariant linear map.
Proof.
$(i)\Rightarrow(ii)$ Suppose $e^{t\mathcal{L}}$ is mixed Weyl-unitary channel for every $t\geq 0$. Then for each $t\geq 0$, there exist finitely many Weyl unitaries $\{W_{j}(t)\}_{j=1}^{n(t)}$ and scalars $\{\lambda_{j}(t)\}_{j=1}^{n(t)}\subseteq[0,1]$ such that $\sum_{j=1}^{n(t)}\lambda_{j}(t)=1$ and $e^{t\mathcal{L}}=\sum_{j=1}^{n(t)}\lambda_{j}(t)\mathrm{Ad}_{W_{j}(t)}$. Consequently,
$\frac{e^{t\mathcal{L}}-\operatorname{\textsf{id}}}{t}=\sum_{j=1}^{n(t)}\lambda_{j}(t)(\mathrm{Ad}_{W_{j}(t)}-\operatorname{\textsf{id}})\in\text{cone}\{\mathrm{Ad}_{W}-\operatorname{\textsf{id}}:W\text{ is Weyl unitary}\},$
Since there are only finitely many Weyl unitaries, so the generating set is finite and hence $\text{cone}\{\mathrm{Ad}_{W}-\operatorname{\textsf{id}}:W\text{ is Weyl unitary}\}$ is closed. Thus, the map $\mathcal{L}=\lim_{t}\frac{e^{t\mathcal{L}}-\operatorname{\textsf{id}}}{t}$ is in the cone generated by $\{\mathrm{Ad}_{W}-\operatorname{\textsf{id}}:W\text{ is Weyl unitary}\}$.
$(ii)\Rightarrow(ii)$ Note that $\mathrm{Ad}_{W}$ is Weyl-covariant map for any Weyl-unitary $W$. Consequently, any map $\Psi$ in the cone generated by $\{\mathrm{Ad}_{W}-\operatorname{\textsf{id}}:W\text{ is Weyl unitary}\}$ is Weyl-covariant. In particular, $\mathcal{L}$ is Weyl-covariant.
$(iii)\Rightarrow(i)$ Assume that $\mathcal{L}$ is a Weyl-covariant linear map. Observe that if $\mathcal{L}$ is Weyl-covariant, then $\mathcal{L}^{k}$ is also Weyl-covariant. Therefore, for every $a,b\in\mathbb{Z}_{d}$ and $X\in\mathbb{M}_{d}$ we have
$e^{t\mathcal{L}}(W_{a,b}XW_{a,b}^{*})$ $=\lim_{n\to\infty}\sum_{k=1}^{n}\frac{t^{k}\mathcal{L}^{k}(W_{a,b}XW_{a,b}^{*})}{k!}$
$=\lim_{n\to\infty}\sum_{k=1}^{n}W_{a,b}\frac{t^{k}\mathcal{L}^{k}X)}{k!}W_{a,b}^{*}$
$=W_{a,b}\big(\lim_{n\to\infty}\sum_{k=1}^{n}\frac{t^{k}\mathcal{L}^{k}X)}{k!}\big)W_{a,b}^{*}$
$=W_{a,b}e^{t\mathcal{L}}(X)W_{a,b}^{*}.$
Thus $e^{t\mathcal{L}}$ is Weyl-covairant for every $t\geq 0$ and by Proposition 6.2 it follows $e^{t\mathcal{L}}$ is mixed Weyl-unitary for every $t\geq 0$. ∎
Note that Weyl unitaries along with $d$-th roots of unity form a finite subgroup of the group of unitaries. Most results of this Section extend to such a general context.
Definition 6.5.
Let $G$ be a subgroup of $\mathbb{U}(d)$. An unital channel $\Phi$ on $\mathbb{M}_{d}$ is said to be $G$-mixed unitary if if there exist $\{U_{j}\}_{j=1}^{\ell}\subseteq G$ and $\{\lambda_{j}\}_{j=1}^{\ell}\subseteq[0,1]$ for some $l\in\mathbb{N}$ such that $\sum_{j=1}^{\ell}\lambda_{j}=1$ and
$\Phi(X)=\sum_{i=1}^{\ell}\lambda_{i}U_{i}^{*}XU_{i},\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \forall X\in\mathbb{M}_{d}.$ (6.2)
Note that any $G$-mixed unitary channel is also a mixed unitary channel. Therefore, the map $\Phi$ given in the Example 2.5 is an example of a non $G$-mixed unitary map on $\mathbb{M}_{3}$.
Denote $\mathscr{MU}_{G}(d)$ be the set of all $G$-mixed unitary channels on $\mathbb{M}_{d}$. Consider $\mathscr{MU}_{G}(d)$ as a subset of the real vector space $\mathscr{H}(d)$.
Lemma 6.6.
For any compact subgroup $G$ of $\mathbb{U}(d)$, we have the following;
  1. (i)
    $\mathscr{MU}_{G}(d)$ is a compact convex set and is closed under compositions;
  2. (ii)
    $\mathscr{MU}_{G}(d)^{\circ}=\{\Gamma\in\mathscr{H}(d):\langle\Gamma,\mathrm{Ad}_{U}\rangle\geq 0\penalty 10000\ \penalty 10000\ \forall U\in G\}$;
  3. (iii)
    Suppose $\Phi\in\mathscr{H}(d)$ is unital and trace-preserving. Then $\Phi\in\mathscr{MU}_{G}(d)$ if and only if $\langle\Gamma,\Phi\rangle\geq 0$ for all unital and trace-preserving $\Gamma\in\mathscr{MU}_{G}(d)^{\circ}$;
  4. (iv)
    The interior of the set $\mathscr{MU}_{G}(d)$ is empty.
Proof.
Proof is similar to the Lemma 2.7. ∎
Denote $\mathscr{D}_{G}(d):=\{\Gamma\in\mathscr{W}(d):\langle\operatorname{\textsf{id}},\Gamma\rangle=0,\langle\mathrm{Ad}_{U},\Gamma\rangle\geq 0,\penalty 10000\ \penalty 10000\ \forall U\in G\}$.
Lemma 6.7.
Suppose $\mathcal{L}\in\mathscr{H}(d)$ with $\langle\Gamma,\mathcal{L}\rangle\geq 0$ for all $\Gamma\in\mathscr{D}_{G}(d)$. Then, for every $r>0$ there exists $\eta>0$ such that if $\Gamma\in\mathcal{C}(d)^{\circ}\cap\mathscr{W}_{1}(d)$ satisfies $\langle\Gamma,\operatorname{\textsf{id}}\rangle<\eta$ then $\langle\Gamma,\mathcal{L}\rangle\geq-r$.
Proof.
Proof similar to Lemma 5.1. ∎
Theorem 6.8.
Let $\mathcal{L}\in\mathscr{H}(d)$ be such that $\mathcal{L}(I_{d})=0$ and $\operatorname{\textsf{tr}}(\mathcal{L}(X))=0,\forall X\in\mathbb{M}_{d}$. Then, the following are equivalent:
  1. (i)
    $e^{t\mathcal{L}}$ is $G$-mixed unitary for all $t\geq 0$;
  2. (ii)
    The generator $\mathcal{L}$ is in the closed cone generated by $\{\mathrm{Ad}_{U}-\operatorname{\textsf{id}}:U\in G\}$;
  3. (iii)
    $\langle\Gamma,\mathcal{L}\rangle\geq 0$ for all $\Gamma\in\mathscr{D}_{G}(d)$.
Proof.
Follows in the same lines of Theorem 5.2. ∎
Note that if the group is finite, in part (ii) of this theorem it suffices to take the non-negative linear combinations of maps of the form $\{\mathrm{Ad}_{U}-\operatorname{\textsf{id}}:U\in G\}$ as there is no need to take the closure.
Acknowledgements
Bhat gratefully acknowledges funding from ANRF (India) through J C Bose Fellowship No. JBR/2021/000024. Devendra acknowledges funding from the Indian Institute of Technology Bombay through the Institute Postdoctoral Fellowship and also thanks the Indian Statistical Institute Bangalore for its kind hospitality during his visits.